### v1

# Complexity Bounds for Smooth Convex Multiobjective Optimization

Abstract
We study the oracle complexity of finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization with $m$ objectives. The progress metric is the Pareto stationarity gap $\mathcal{G}(x)$ (the norm of an optimal convex combination of gradients). Our contributions are fourfold. (i) For strongly convex objectives, any span first-order method (iterates lie in the span of past gradients) exhibits linear convergence no faster than $\exp(-\Theta(T/\sqrt{\kappa}))$ after $T$ oracle calls, where $\kappa$ is the condition number, implying $\Theta(\sqrt{\kappa}\log(1/\varepsilon))$ iterations; this matches classical accelerated upper bounds. (ii) For convex problems and oblivious one-step methods (a fixed scalarization with pre-scheduled step sizes), we prove a lower bound of order $1/T$ on the best gradient norm among the first $T$ iterates. (iii) Although accelerated gradient descent is outside this restricted class, it is an oblivious span method and attains the same $1/T$ upper rate on a fixed scalarization. (iv) For convex problems and general span methods with adaptive scalarizations, we establish a universal lower bound of order $1/T^{2}$ on the gradient norm of the final iterate after $T$ steps, highlighting a gap between known upper bounds and worst-case guarantees. All bounds hold on non-degenerate instances with distinct objectives and non-singleton Pareto fronts; rates are stated up to universal constants and natural problem scaling.
Keywords Multiobjective optimization; Oracle complexity; Pareto stationarity; First-order span methods; Adaptive scalarization
1 Introduction
Multiobjective optimization (MOO) provides a mathematical framework for decision-making problems involving multiple, often conflicting, performance criteria. Given a vector-valued objective function $F(x)=(f_{1}(x),\dots,f_{m}(x))$, the goal is to find a point $x$ that offers the best possible trade-off among the individual objectives $f_{i}$. Since a single point that simultaneously minimizes all objectives typically does not exist, the central solution concept is that of Pareto optimality. A point is Pareto optimal if no other point can improve one objective without degrading at least one other.
First-order iterative methods have become the tool of choice for solving large-scale MOO problems, particularly in machine learning [29], and they are increasingly deployed in engineering and finance — for example, distributed-gradient schemes for multi-objective AC optimal power flow [22], adjoint-based gradient methods for reservoir well-control under risk and uncertainty [19], and gradient-descent formulations for multi-objective portfolio selection [26]. These methods rely on an oracle that returns function values and gradients. A key challenge is to define a tractable computational goal. While computing an exact Pareto-optimal point is often impractical, a widely adopted necessary condition is Pareto criticality [13], which is equivalent to the Pareto stationarity notion adopted here. A point $x$ is Pareto stationary if the convex hull of its gradients contains the origin. The proximity to this condition is measured by the Pareto stationarity gap $\mathcal{G}(x)$ (see Definition 3.5 in Section 3). Finding a point $x$ with $\mathcal{G}(x)\leq\varepsilon$ is a standard objective for many contemporary MOO algorithms [10].
Although numerous algorithms have been proposed with convergence-rate guarantees (upper bounds), the fundamental limits of what is achievable have remained largely unexplored. Without oracle lower bounds, we cannot know if existing algorithms are optimal. In single-objective optimization, this story is complete: the seminal work of Nemirovski and Yudin [23] established lower bounds that were later matched by accelerated methods [24], proving their optimality. For MOO, however, a corresponding theory of oracle complexity has been absent.
Why scalarization does not trivialize MOO.
It may appear that multiobjective complexity reduces immediately to the single-objective case by fixing a scalarization $f_{\lambda}=\sum_{i}\lambda_{i}f_{i}$ and applying known convergence results. Indeed, all classical rates for smooth convex or strongly convex minimization (e.g., gradient descent $O(LR/\sqrt{T})$, AGD $O(LR/T)$, and linear convergence under strong convexity) apply verbatim to any fixed scalarization. However, the multiobjective goal is different: the central complexity measure is the Pareto stationarity gap $\mathcal{G}(x)$ which requires considering all convex combinations of gradients simultaneously. A point minimizing a fixed $f_{\lambda}$ need not minimize $\mathcal{G}$, and the algorithm classes we study (oblivious one-step vs. span/adaptive methods) differ in their ability to exploit information across objectives. Consequently, while scalarization provides useful upper bounds, the oracle complexity of Pareto stationarity requires a dedicated analysis, which is the focus of this paper.
Contributions.
This paper closes this fundamental gap by providing a nuanced information-based complexity analysis that distinguishes between different classes of first-order algorithms. Our main contributions are as follows:
  • Tight Bounds for Strongly Convex MOO: We establish a tight linear convergence lower bound. Any span first-order method can be forced to have a stationarity gap of at least $\mu R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T}$, implying an iteration complexity of $\Omega(\sqrt{\kappa}\log(\mu R/\varepsilon))$. This is matched by accelerated gradient methods on a fixed scalarization.
  • Lower Bound for Oblivious One-Step Convex MOO: For the class of oblivious one-step gradient methods—which includes gradient descent with a pre-scheduled step-size sequence—we prove an $\Omega(LR/(T+1))$ lower bound on the (min-iterate) stationarity gap. Accelerated gradient descent (AGD), which is not in this class but is an oblivious span method, achieves an $\mathcal{O}(LR/(T+1))$ upper rate on a fixed scalarization. Thus the $\Omega(1/T)$ rate is matched by oblivious span methods, while within the oblivious one-step class the best known general upper bound is $\mathcal{O}(LR/\sqrt{T})$.
  • Universal Lower Bound for Adaptive Methods: For fully adaptive span first-order methods, we establish a universal last-iterate lower bound of $\Omega(LR/(T+1)^{2})$. We highlight the remaining gap between this and the $\mathcal{O}(1/T)$ upper bounds of optimal methods as a key open problem.
  • Robustness of Constructions: We show our lower bounds are not artifacts of degenerate problems. They hold for instances with distinct, conflicting objectives and non-singleton Pareto fronts.
Our methodology refines the classical resisting oracle framework [23] by carefully analyzing the polynomial representations of different algorithm classes. These results provide a definitive characterization of the oracle complexity for a central task in modern multiobjective optimization.
Organization.
After surveying the literature in the Related Work section, Section 3 fixes notation, introduces the Pareto stationarity gap $\mathcal{G}(\cdot)$ and its equivalence to Pareto criticality, and formalizes the oracle model and algorithm classes (span methods vs. oblivious one-step), including the Krylov/span polynomial representation. Section 4 develops our lower bounds: a tight linear-rate bound in the strongly convex case (Theorem 4.1), an $\Omega(1/T)$ min-iterate bound for oblivious one-step methods (Theorem 4.3), and a universal $\Omega(1/T^{2})$ last-iterate bound for fully adaptive span methods (Theorem 4.4), together with an explicit non-degenerate construction showing robustness of the instances (Section 4.3). Section 5 provides matching upper bounds via AGD on fixed scalarizations and compares rates against the lower bounds. Section 6 concludes with implications and open questions. The appendices contain the polynomial extremal derivations (Chebyshev and Markov) and auxiliary proofs used throughout.
2 Related Work
Oracle complexity and first-order limits.
The information-based complexity program for single-objective convex optimization established sharp first-order lower bounds together with matching optimal methods [23, 24]. In [23], these lower bounds are derived by reducing the worst-case error to an extremal polynomial problem, solved exactly by (scaled) Chebyshev polynomials; see [28] for the underlying Chebyshev minimax and related inequalities. Recent analyses include Performance Estimation Problem (PEP) frameworks that deliver tight worst-case guarantees for many first-order methods [8]. Beyond the classical oracle model, a line of work formalizes oblivious and structured first-order algorithms: Arjevani and Shamir [3] introduced the oblivious framework with dimension-free lower bounds for $L$-smooth convex and $L$-smooth $\mu$-strongly convex problems; Arjevani, Shalev-Shwartz, and Shamir [1] developed the $p$-SCLI framework (covering p=1p{=}1 one-step methods) with matching bounds; and Arjevani and Shamir [2] extended dimension-free iteration lower bounds to finite-sum settings, covering variance-reduced families. Complementary results for randomized first-order algorithms appear in Woodworth and Srebro [32]. A comprehensive and recent treatment of finite-sum lower bounds is given by Han, Xie, and Zhang [18].
Multiobjective optimization algorithms.
Classical foundations and broad context are covered in the monograph by Miettinen [21] and the survey of Marler and Arora [20]. For descent-type methods, the steepest–descent family beginning with [13] and including cone-ordered variants and refined convergence analyses [9, 7] provides global convergence to Pareto-stationary points. Complexity aspects for gradient descent in smooth MOO are analyzed in [11]. Universal nonmonotone line-search frameworks for non-convex MOO with convex constraints have also been proposed [27]. When progress is tracked through standard merit functions or stationarity surrogates, the resulting upper bounds are typically sublinear (e.g., $\mathcal{O}(1/k)$) in smooth convex settings [31, 16]. Second-order frameworks — Newton-type and SQP — offer globalization with fast local behavior: globally convergent Newton methods for MOO with fast (often superlinear) local rates under standard regularity are developed in [17], while SQP-based schemes for constrained problems are given in [12]. Conditional-gradient (Frank–Wolfe) methods have been developed for vector/multiobjective optimization with global convergence to weakly efficient (Pareto-stationary) solutions under cone-convexity assumptions [6]. Away-step and related variants can accelerate convergence and, under additional geometric or curvature assumptions, deliver faster (sometimes linear) rates [15, 16], with adaptive generalized conditional-gradient methods extending this line [14]. Accelerated first-order schemes tailored to the multiobjective setting continue to mature: an accelerated proximal-gradient method with $\mathcal{O}(1/k^{2})$ decrease for an appropriate merit measure is established in [30], and recent work clarifies monotonicity properties of multiobjective APG [25]. Complementary practical gradient-based schemes and stationarity criteria, such as multiple-gradient descent, are surveyed in [10], and links to multi-task learning further motivate stationarity-based goals and algorithm design [29].
Our position.
The works above emphasize algorithm design and, in the scalar case, tight lower bounds under various oracle or structural models. To our knowledge, a systematic oracle-complexity theory for Pareto stationarity in smooth MOO — distinguishing oblivious one-step methods from general span first-order methods and providing robust non-degenerate lower-bound instances — has not been previously developed. The present paper fills this gap and connects multiobjective first-order complexity to the classical scalar theory through polynomial extremal tools.
3 Preliminaries
3.1 Notation
All vectors lie in $\mathbb{R}^{d}$ and matrices in $\mathbb{R}^{d\times d}$. The Euclidean inner product and norm are $\langle x,y\rangle$ and $\|x\|$; for a matrix $A$, $\|A\|_{\mathrm{op}}$ is the operator norm, $\sigma(A)$ its spectrum, $A\succeq 0$ (resp. $A\succ 0$) means positive (semi)definite, and $I$ is the identity; $(\cdot)^{\top}$ denotes transpose. For a set $S$, $\operatorname{conv}(S)$ is its convex hull, $\operatorname{span}(S)$ its linear span, and $\operatorname{dist}(x,S):=\inf_{y\in S}\|x-y\|$ its distance. For a matrix $A$, $\operatorname{range}(A)$ and $\ker(A)$ denote its range and kernel. We write $\mathbb{N}_{0}:=\{0,1,2,\ldots\}$ and $\mathbb{N}:=\{1,2,\ldots\}$.
For multiobjective problems with $m\geq 2$ objectives $F(x)=(f_{1}(x),\ldots,f_{m}(x))$, the unit simplex is $\Delta_{m}:=\{\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\}$. For $\lambda\in\Delta_{m}$, the scalarization is $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ and $x^{\star}_{\lambda}\in\arg\min_{x}f_{\lambda}(x)$ denotes an arbitrary minimizer. The Pareto set is denoted by $\mathcal{P}$. We use $R:=\operatorname{dist}(x(0),\mathcal{P})$ and $R_{\lambda}:=\|x(0)-x^{\star}_{\lambda}\|$ for initialization radii. Iterates are $x(t)$ for $t\in\mathbb{N}_{0}$; $T\in\mathbb{N}$ is the number of oracle calls/iterations. The condition number is $\kappa:=L/\mu$ when strong convexity applies; we occasionally write $\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}$ in spectral formulas. For quadratics $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, we use the spectral variable $\zeta$ and real polynomials $p_{t}$ (with $\deg p_{t}\leq t$, $p_{t}(0)=1$) to express errors via $x(t)-x^{\star}=p_{t}(H)\bigl{(}x(0)-x^{\star}\bigr{)}$; Chebyshev polynomials of the first kind are denoted $T_{t}(\cdot)$. Standard Landau symbols $O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ are used; $\log$ denotes the natural logarithm. The vectors $(e_{i})_{i=1}^{m}$ are the canonical basis of $\mathbb{R}^{m}$.
We consider unconstrained optimization problems in $\mathbb{R}^{d}$. A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth if its gradient is $L$-Lipschitz continuous. It is $\mu$-strongly convex if $f(y)\geq f(x)+\left\langle\nabla f(x),y-x\right\rangle+\frac{\mu}{2}\left\lVert y-x\right\rVert^{2}$ for all $x,y$.
3.2 Problem formulation
We consider unconstrained, continuously differentiable multiobjective problems
$\min_{x\in\mathbb{R}^{d}}F(x):=(f_{1}(x),\ldots,f_{m}(x)).$ (1)
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex and $L$-smooth.
3.3 Pareto optimality and first-order conditions
We now review the standard optimality notions for (possibly non-convex) MOO problems (see [21]) and clarify their relationship with the Pareto stationarity gap used throughout the paper.
Definition 3.1 (Dominance).
Given $x,y\in\mathbb{R}^{d}$, we say that $y$ dominates $x$ if
$f_{i}(y)\leq f_{i}(x)\ \ \text{for all }i=1,\dots,m,\quad\text{and}\quad f_{j}(y)<f_{j}(x)\ \ \text{for at least one }j.$
Based on this, a point is considered optimal if no other point in the domain dominates it.
Definition 3.2 (Pareto optimality (strong)).
A point $x^{\star}\in\mathbb{R}^{d}$ is Pareto optimal if there is no $y\in\mathbb{R}^{d}$ that dominates $x^{\star}$; equivalently, there is no $y$ with
$f_{i}(y)\leq f_{i}(x^{\star})\ \ \text{for all }i,\quad\text{and}\quad f_{j}(y)<f_{j}(x^{\star})\ \ \text{for some }j.$
A related, slightly weaker condition is often useful for theoretical analysis and algorithm design:
Definition 3.3 (Weak Pareto optimality).
A point $x^{\star}\in\mathbb{R}^{d}$ is weakly Pareto optimal if there is no $y\in\mathbb{R}^{d}$ such that
$f_{i}(y)<f_{i}(x^{\star})\ \ \text{for all }i=1,\dots,m.$
To develop iterative, gradient-based methods, we need a computable, first-order necessary condition for optimality. This leads to the concept of Pareto criticality.
Definition 3.4 (Pareto criticality).
Following [13], a point $x\in\mathbb{R}^{d}$ is called Pareto critical if there is no common strict descent direction, i.e.,
$\not\exists\,v\in\mathbb{R}^{d}\quad\text{such that}\quad\nabla f_{i}(x)^{\top}v<0\quad\text{for all }i=1,\dots,m.$
Equivalently, with the Jacobian $JF(x)\in\mathbb{R}^{m\times d}$ whose $i$-th row is $\nabla f_{i}(x)^{\top}$, this reads
$\operatorname{range}(JF(x))\cap(-\mathbb{R}_{++})^{m}=\varnothing,$
where $\mathbb{R}_{++}$ is the set of (strictly) positive real numbers.
This geometric condition is equivalent to the convex hull of the gradients containing the origin. The proximity to this state is quantified by the following gap function.
Definition 3.5 (Pareto stationarity gap).
The Pareto stationarity gap at $x\in\mathbb{R}^{d}$ is
$\mathcal{G}(x)\;:=\;\min_{\lambda\in\Delta^{m}}\Big{\|}\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)\Big{\|},\qquad\Delta^{m}:=\big{\{}\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\big{\}}.$
We call $x$ $\varepsilon$-Pareto stationary if $\mathcal{G}(x)\leq\varepsilon$; the case $\varepsilon=0$ corresponds to Pareto stationarity.
Equivalence of stationarity notions.
The next lemma makes precise that Pareto criticality is equivalent to Pareto stationarity $\mathcal{G}(x)=0$.
Lemma 3.6 (No common descent $\Longleftrightarrow$ convex-hull stationarity).
For $C^{1}$ objectives $(f_{i})$, the following are equivalent at a point $x$:
  1. (a)
    (Pareto criticality) There is no $v\in\mathbb{R}^{d}$ with $\nabla f_{i}(x)^{\top}v<0$ for all $i$.
  2. (b)
    (Convex-hull stationarity) There exists $\lambda\in\Delta^{m}$ such that $\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)=0$ (equivalently, $\mathcal{G}(x)=0$).
Proof.
Apply Gordan’s theorem of the alternative to $A=JF(x)$: exactly one of $Av<0$ or $A^{\top}y=0$ with $y\geq 0$, $y\neq 0$ holds. The negation of the former yields a nonzero $y\geq 0$ with $JF(x)^{\top}y=0$; normalizing $y$ to the simplex gives (b). Conversely, (b) rules out any $v$ with $JF(x)v<0$ because $\sum_{i}\lambda_{i}\,\nabla f_{i}(x)^{\top}v$ would be a convex combination of strictly negative numbers. ∎
Necessary and sufficient conditions.
We now record the basic relationships between optimality and stationarity in the unconstrained setting.
Proposition 3.7 (General $C^{1}$ case: necessity).
If $x^{\star}$ is weakly Pareto optimal (for not-necessarily convex $C^{1}$ objectives), then $x^{\star}$ is Pareto stationary; equivalently $\mathcal{G}(x^{\star})=0$.
Proof.
At a weakly Pareto optimal point there is no direction that strictly decreases all objectives. By Lemma 3.6, this is equivalent to the existence of $\lambda\in\Delta^{m}$ with $\sum_{i}\lambda_{i}\,\nabla f_{i}(x^{\star})=0$, hence $\mathcal{G}(x^{\star})=0$. ∎
Proposition 3.8 (Convex $C^{1}$ case: characterization of weak Pareto optima).
Suppose each $f_{i}$ is convex and $C^{1}$. Then, for $x^{\star}\in\mathbb{R}^{d}$, the following are equivalent:
  1. (i)
    $x^{\star}$ is weakly Pareto optimal;
  2. (ii)
    $x^{\star}$ is Pareto stationary, i.e., $\mathcal{G}(x^{\star})=0$;
  3. (iii)
    there exists $\lambda\in\Delta^{m}$ such that $x^{\star}\in\arg\min_{x}f_{\lambda}(x)$ with $f_{\lambda}:=\sum_{i=1}^{m}\lambda_{i}f_{i}$.
Proof.
[…]
Remark 3.9.
Under convexity, Proposition 3.8 shows that $\mathcal{G}(\cdot)$ is an exact first-order optimality measure for weak Pareto optimality: $\mathcal{G}(x^{\star})=0$ if and only if $x^{\star}$ is weakly Pareto optimal. In the general (nonconvex) $C^{1}$ case, $\mathcal{G}(x^{\star})=0$ remains a necessary condition but is not sufficient, as usual for first-order stationarity.
3.4 Oracle Model and Algorithm Classes
We formally define the classes of first-order algorithms considered in this work. The broadest class contains nearly all modern iterative methods.
Definition 3.10 (Span First-Order Methods).
An algorithm is a span first-order method if its iterates satisfy
$x^{(t)}\;\in\;x^{(0)}+\mathrm{span}\{\nabla f_{i}(x^{(j)}):\ i=1,\dots,m,\ j=0,\dots,t-1\}.$
This general class allows for adaptive step sizes and momentum terms that depend on the entire history of observed gradients.
For our sharpest convex lower bound, we consider a more restricted, non-adaptive class.
Definition 3.11 (Oblivious One-Step Gradient Methods).
Fix $\lambda\in\Delta^{m}$ and define $f_{\lambda}(x)=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$. An oblivious one-step gradient method is an algorithm that generates iterates by
$x^{(t+1)}=x^{(t)}-\alpha_{t}\,\nabla f_{\lambda}(x^{(t)}),\qquad t=0,1,2,\dots,$
with pre-scheduled step sizes $\{\alpha_{t}\}$ satisfying $0\leq\alpha_{t}\leq 1/L$ for all $t$ that may depend only on known problem parameters (e.g., $L$, $\mu$), but not on the oracle’s feedback. For a quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, the error satisfies $x^{(t)}-x^{\star}=p_{t}(H)(x^{(0)}-x^{\star})$ with
$p_{t}(\zeta)\;=\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\zeta\bigr{)},\qquad p_{t}(0)=1.$
Remark 3.12.
The oblivious one-step class models standard gradient descent with a fixed step schedule. It does not include methods with momentum, such as Nesterov’s accelerated gradient method (AGD) or Polyak’s heavy-ball method.
Lemma 3.13 (Krylov Representation for Span Methods on Quadratics).
Let $g:\mathbb{R}^{d}\to\mathbb{R}$ be the quadratic
$g(x)=\tfrac{1}{2}\,x^{\top}Hx-b^{\top}x,$
where $H\in\mathbb{R}^{d\times d}$ is symmetric and $b\in\operatorname{range}(H)$, so that $g$ attains its minimum and there exists $x^{\star}\in\arg\min g$ with $Hx^{\star}=b$. Consider any iterative algorithm whose iterates $\{x^{(t)}\}_{t\geq 0}$ satisfy the span property
$x^{(t)}\in x^{(0)}+\operatorname{span}\{\nabla g(x^{(0)}),\ldots,\nabla g(x^{(t-1)})\}\quad\text{for all }t\geq 1.$ (2)
Then, for every $t\geq 0$, there exists a real polynomial $p_{t}$ of degree at most $t$ with $p_{t}(0)=1$ such that
$x^{(t)}-x^{\star}\;=\;p_{t}(H)\,\bigl{(}x^{(0)}-x^{\star}\bigr{)}.$
Moreover, if the method is an oblivious one-step gradient method on $g$, i.e.,
$x^{(t+1)}=x^{(t)}-\alpha_{t}\,\nabla g(x^{(t)})\quad(t\geq 0)$
for some pre-scheduled stepsizes $\{\alpha_{t}\}_{t\geq 0}$, then
$p_{t}(\zeta)=\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)}\qquad(t\geq 1),\qquad p_{0}\equiv 1.$
The same conclusions hold when $g$ is any fixed scalarization of a quadratic multiobjective instance.
Proof.
Fix any minimizer $x^{\star}$ with $Hx^{\star}=b$; such a point exists by the assumption $b\in\operatorname{range}(H)$. For every $x$, we then have
$\nabla g(x)=Hx-b=H\,(x-x^{\star}).$ (3)
Let $e^{(t)}:=x^{(t)}-x^{\star}$ and $e^{(0)}:=x^{(0)}-x^{\star}$. We prove by induction on $t$ that
$e^{(t)}\;=\;p_{t}(H)\,e^{(0)}\quad\text{for some polynomial }p_{t}\text{ with }\deg p_{t}\leq t\text{ and }p_{t}(0)=1.$ (4)
Base case $t=0$. Trivially, $e^{(0)}=I\,e^{(0)}$, so (4) holds with $p_{0}\equiv 1$.
Inductive step. Assume (4) holds for all indices up to $t$. By the span property (2), there exist (possibly data-dependent) scalars $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ such that
$x^{(t+1)}\;=\;x^{(0)}+\sum_{j=0}^{t}\beta_{j}^{(t)}\,\nabla g\bigl{(}x^{(j)}\bigr{)}.$
Subtracting $x^{\star}$ and using (3) gives
$e^{(t+1)}\;=\;e^{(0)}+\sum_{j=0}^{t}\beta_{j}^{(t)}\,H\,e^{(j)}.$ (5)
By the induction hypothesis, for each $0\leq j\leq t$ there exists a polynomial $p_{j}$ with $\deg p_{j}\leq j$ and $p_{j}(0)=1$ such that $e^{(j)}=p_{j}(H)\,e^{(0)}$. Substituting into (5) yields
$e^{(t+1)}\;=\;\Bigl{(}I+H\,\sum_{j=0}^{t}\beta_{j}^{(t)}\,p_{j}(H)\Bigr{)}\,e^{(0)}.$
Define the polynomial
$q_{t}(\zeta)\;:=\;\sum_{j=0}^{t}\beta_{j}^{(t)}\,p_{j}(\zeta),\qquad\text{so that}\qquad\deg q_{t}\leq\max_{0\leq j\leq t}\deg p_{j}\leq t,$
and set
$p_{t+1}(\zeta)\;:=\;1+\zeta\,q_{t}(\zeta).$
Then $\deg p_{t+1}\leq t+1$, $p_{t+1}(0)=1$, and
$e^{(t+1)}\;=\;p_{t+1}(H)\,e^{(0)}.$
This completes the induction and proves (4) for all $t\geq 0$.
It remains to establish the stated product form for oblivious one-step methods. In that case,
$e^{(t+1)}\;=\;x^{(t)}-x^{\star}-\alpha_{t}\,\nabla g(x^{(t)})\;=\;\bigl{(}I-\alpha_{t}H\bigr{)}\,e^{(t)}\qquad(t\geq 0),$
using (3). Iterating this linear recurrence and recalling $e^{(0)}=p_{0}(H)\,e^{(0)}$ with $p_{0}\equiv 1$ gives
$e^{(t)}\;=\;\Bigl{(}\prod_{k=0}^{t-1}\bigl{(}I-\alpha_{k}H\bigr{)}\Bigr{)}\,e^{(0)}\;=\;p_{t}(H)\,e^{(0)},$
with $p_{t}(\zeta)=\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)$, which has degree $t$ and satisfies $p_{t}(0)=1$.
Independence of the choice of minimizer. If $H$ is singular, the minimizer may not be unique: any $x^{\star}$ with $Hx^{\star}=b$ differs from another minimizer by a vector in $\ker(H)$. Let $x^{\star}{}^{\prime}$ be another minimizer and write $x^{\star}{}^{\prime}=x^{\star}+z$ with $Hz=0$. Then $e^{\prime(0)}=x^{(0)}-x^{\star}{}^{\prime}=e^{(0)}-z$ and, by the proven representation,
$x^{(t)}-x^{\star}{}^{\prime}\;=\;p_{t}(H)\,e^{\prime(0)}\;=\;p_{t}(H)\,e^{(0)}-p_{t}(0)\,z.$
Since $x^{(t)}-x^{\star}{}^{\prime}=(x^{(t)}-x^{\star})-z$, consistency for all $z\in\ker(H)$ requires and is ensured by $p_{t}(0)=1$. Thus the representation is well defined independently of the chosen minimizer. ∎
4 Oracle Complexity Lower Bounds
We present our lower bounds by constructing "resisting oracle" instances. In all constructions, we define the initial distance to the Pareto set as $R:=\mathrm{dist}(x^{(0)},\mathcal{P})$. For any specific scalarization $\lambda\in\Delta^{m}$, the distance to its unique minimizer $x_{\lambda}^{\star}$ satisfies $R_{\lambda}:=\|x^{(0)}-x^{\star}_{\lambda}\|\geq R$. Our hard instances are constructed such that we can choose a $\lambda$ for which $R_{\lambda}=R$, making the bounds tight.
4.1 The Strongly Convex Case: A Tight Linear Rate
We begin by establishing a lower bound on the convergence rate for strongly convex multiobjective optimization. Our analysis considers the broad class of span first-order methods and demonstrates that any such algorithm is subject to a linear convergence rate fundamentally limited by the problem’s condition number. The following theorem formalizes this result.
Theorem 4.1 (Strongly Convex Lower Bound).
Fix $L\geq\mu>0$, $T\in\mathbb{N}$, $\kappa:=L/\mu$, and $R>0$. Consider any span first-order method that makes $T$ first-order oracle calls when applied to a quadratic objective. Then there exists a symmetric positive definite matrix $H\in\mathbb{R}^{d\times d}$ with spectrum $\sigma(H)\subset[\mu,L]$ (for some $d\geq T+1$), a vector $b\in\mathbb{R}^{d}$, and an initialization $x^{(0)}\in\mathbb{R}^{d}$ such that for the quadratic
$g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x,\qquad x^{\star}=H^{-1}b,\qquad\|x^{(0)}-x^{\star}\|=R,$
the $T$-th iterate produced by the method satisfies
$\bigl{\|}\nabla g\bigl{(}x^{(T)}\bigr{)}\bigr{\|}\;\geq\;\mu\,R\;\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
In particular, any such method has worst-case iteration complexity $\Omega\!\bigl{(}\sqrt{\kappa}\log(1/\varepsilon)\bigr{)}$ to reduce the gradient norm below $\varepsilon$.
Proof.
Let $g(x)=\frac{1}{2}x^{\top}Hx-b^{\top}x$ with $H\succ 0$. For any iterate sequence $\{x^{(t)}\}$ produced by a span first-order method on a quadratic, the error admits the polynomial representation (Lemma 3.13):
$x^{(t)}-x^{\star}\;=\;p_{t}(H)\,\bigl{(}x^{(0)}-x^{\star}\bigr{)},\qquad\deg p_{t}\leq t,\quad p_{t}(0)=1.$ (6)
Set $e^{(0)}:=x^{(0)}-x^{\star}$ with $\|e^{(0)}\|=R$. Since $H\succeq\mu I$, we have for any vector $v$ that $\|Hv\|\geq\mu\|v\|$, hence
$\bigl{\|}\nabla g(x^{(T)})\bigr{\|}=\bigl{\|}H\,p_{T}(H)\,e^{(0)}\bigr{\|}\;\geq\;\mu\,\bigl{\|}p_{T}(H)\,e^{(0)}\bigr{\|}.$ (7)
We now take worst case over the instance $(H,e^{(0)})$ with $\sigma(H)\subset[\mu,L]$ and $\|e^{(0)}\|=R$. For any fixed degree-$T$ polynomial $p_{T}$ with $p_{T}(0)=1$,
$\sup_{\begin{subarray}{c}\sigma(H)\subset[\mu,L]\\ \|e^{(0)}\|=R\end{subarray}}\bigl{\|}p_{T}(H)\,e^{(0)}\bigr{\|}\;=\;R\;\sup_{\zeta\in[\mu,L]}|p_{T}(\zeta)|,$
because for $H$ symmetric the operator norm $\|p_{T}(H)\|$ equals $\max_{\zeta\in\sigma(H)}|p_{T}(\zeta)|$, and we may align $e^{(0)}$ with an eigenvector where the maximum is attained. Therefore, combining with (7) and then minimizing over the (algorithm-induced) polynomial $p_{T}$ yields
$\sup_{\begin{subarray}{c}\sigma(H)\subset[\mu,L],\,\|e^{(0)}\|=R\end{subarray}}\bigl{\|}\nabla g(x^{(T)})\bigr{\|}\;\geq\;\mu R\cdot\inf_{\begin{subarray}{c}\deg p_{T}\leq T\\ p_{T}(0)=1\end{subarray}}\;\max_{\zeta\in[\mu,L]}|p_{T}(\zeta)|.$ (8)
The minimax problem on the right-hand side is the classical Chebyshev extremal problem on an interval with an off-interval normalization (at $\zeta=0$). The solution is detailed in Appendix A.3. Let
$\xi\;=\;\frac{2\zeta-(L+\mu)}{L-\mu}\in[-1,1],\qquad\xi_{0}\;=\;\frac{-\,(L+\mu)}{L-\mu}\;=\;-\,\frac{\kappa+1}{\kappa-1},$
and define $\tilde{p}_{T}(\xi):=p_{T}(\zeta(\xi))$. Then the extremal value is achieved by the scaled Chebyshev polynomial of the first kind,
$\tilde{p}_{T}^{\star}(\xi)\;=\;\frac{T_{T}(\xi)}{T_{T}(\xi_{0})},\qquad\min_{\deg\tilde{p}_{T}\leq T,\;\tilde{p}_{T}(\xi_{0})=1}\;\max_{\xi\in[-1,1]}|\tilde{p}_{T}(\xi)|\;=\;\frac{1}{|T_{T}(\xi_{0})|}.$
For $|x|>1$ one has $T_{T}(x)=\tfrac{1}{2}\bigl{(}\rho^{T}+\rho^{-T}\bigr{)}$ with $\rho=x+\sqrt{x^{2}-1}$. Substituting $x=|\xi_{0}|=\frac{\kappa+1}{\kappa-1}$ gives
$|T_{T}(\xi_{0})|\;=\;\frac{\rho^{T}+\rho^{-T}}{2},\qquad\rho=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
Thus the extremal value of (8) equals $\mu R\cdot\frac{2}{\rho^{T}+\rho^{-T}}$, proving
$\sup_{\text{instances}}\bigl{\|}\nabla g(x^{(T)})\bigr{\|}\;\geq\;\mu R\cdot\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},$
where the last inequality follows from $\frac{2}{a+a^{-1}}\geq a^{-1}$ for $a\geq 1$. The iteration-complexity statement follows by solving $\mu R\bigl{(}(\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)\bigr{)}^{T}\leq\varepsilon$ for $T$. ∎
Remark 4.2 (Lifting to MOO).
The construction of Theorem 4.5 with $\gamma=\mu$ yields $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$, thus Theorem 4.1 extends directly to MOO.
4.2 The Convex Case: Oblivious vs. Adaptive Methods
We now turn to the smooth convex setting, where strong convexity is no longer assumed. In this regime, the achievable complexity bounds depend crucially on the information available to the algorithm. We therefore draw a distinction between oblivious one-step methods, which employ a fixed scalarization and pre-scheduled step sizes, and fully adaptive span methods. We derive distinct lower bounds for each class, revealing a fundamental separation in their worst-case performance guarantees.
4.2.1 Oblivious One-Step Gradient Methods: An $\Omega(1/T)$ Min-Iterate Bound
For the more restricted class of oblivious one-step gradient methods, we establish a lower bound of order $\Omega(1/T)$. This bound applies to the minimum Pareto stationarity gap achieved over the first $T$ iterates and demonstrates a fundamental performance limit for non-adaptive algorithms that do not incorporate momentum or history-dependent updates.
Theorem 4.3 (Oblivious Convex Lower Bound).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any oblivious one-step gradient method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\frac{LR}{4(T+1)}.$
Proof.
Step 1: Spectral hard instance (diagonal is sufficient). By Lemma 3.13, for an oblivious one-step method on a quadratic with spectrum in $[0,L]$ we have $\nabla g(x^{(t)})=H\,p_{t}(H)\,(x^{(0)}-x^{\star})$ with $p_{t}(\zeta)=\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)$ and $\alpha_{k}\leq 1/L$. For any fixed stepsizes, the worst case over instances is obtained by placing an eigenvalue at a maximizer $\zeta^{\star}\in[0,L]$ of $\zeta\big{|}\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)\big{|}$ and aligning $x^{(0)}-x^{\star}$ with the corresponding eigenvector; in particular, a diagonal (even one-dimensional) $H=\zeta^{\star}$ suffices. Thus the lower bound reduces to the product-form extremal
$\inf_{\{\alpha_{k}\}_{k=0}^{t-1}\subset[0,1/L]}\ \max_{\zeta\in[0,L]}\ \zeta\ \Bigl{|}\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)\Bigr{|}\ \geq\ \frac{L}{4(t+1)},$
proved in Appendix A.1. Taking the minimum over $t\in\{0,\dots,T\}$ is achieved at $t=T$.
Step 2: Lifting to MOO. Using the construction of Theorem 4.5 with this $L$-smooth convex quadratic $g$, we obtain the desired MOO lower bound. ∎
4.2.2 Fully Adaptive Methods: A Universal Last-Iterate Lower Bound and an Open Problem
When the algorithm can be fully adaptive, the error polynomial $p_{t}$ no longer has the restrictive product form. This leads to a weaker, but more general, lower bound based on Markov’s inequality for polynomials.
Theorem 4.4 (Universal Convex Lower Bound).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any span first-order method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the last iterate obeys
$\mathcal{G}(x^{(T)})\;\geq\;\frac{LR}{2\,(T+1)^{2}}.$
Proof.
Step 1: The Hard Scalar Instance. We use the same hard quadratic $g(x)$ as in the proof of Theorem 4.3. For a general span method, the error polynomial $p_{t}$ from Lemma 3.13 can be any polynomial of degree at most $t$ with $p_{t}(0)=1$. The lower bound comes from the extremal problem:
$\inf_{\begin{subarray}{c}\deg p_{t}\leq t\\ p_{t}(0)=1\end{subarray}}\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|\;\geq\;\frac{L}{2(t+1)^{2}}.$
This result is a direct consequence of the Markov brothers’ inequality, which bounds the derivative of a polynomial. The full derivation is provided in Appendix A.2. The minimum over $t\in\{0,\dots,T\}$ is achieved at $t=T$, yielding $\frac{LR}{2(T+1)^{2}}$.
Step 2: Lifting to MOO. The result follows by applying the construction of Theorem 4.5, which translates the scalar gradient norm lower bound into a lower bound on the Pareto stationarity gap. ∎
Iterate selection. Theorem 4.3 is a min-iterate lower bound, whereas Theorem 4.4 is a last-iterate bound; the two statements are therefore not directly comparable a priori.
This $\Omega(1/T^{2})$ bound is consistent with (but weaker than) the $\mathcal{O}(LR/\sqrt{T})$ gradient-norm upper bound of vanilla gradient descent, and leaves a gap to the $\mathcal{O}(LR/T)$ upper bound achieved by AGD.
Open Problem. Does a universal $\Omega(1/T)$ lower bound hold for the Pareto stationarity gap $\min_{t\leq T}\mathcal{G}(x^{(t)})$ for all span first-order methods in smooth convex MOO, or can an adaptive algorithm achieve a rate faster than $\mathcal{O}(1/T)$?
4.3 A Construction for Non-Degenerate Instances
A potential criticism of lower bounds is that they might only hold on degenerate problem instances (e.g., where all $f_{i}$ are identical, or the Pareto set is a single point). We now show that our bounds hold for robust, non-degenerate problems.
Theorem 4.5 (Robustness of Lower Bounds).
The lower bounds stated in Theorems 4.1, 4.3, and 4.4 hold for MOO instances where:
  1. (i)
    The objective functions $\{f_{i}\}_{i=1}^{m}$ are distinct.
  2. (ii)
    The Pareto set $\mathcal{P}$ is a non-singleton convex polytope.
Proof.
Let a span first-order method and $T$ be given. We construct a hard MOO instance. Decompose the space as $\mathbb{R}^{d}=V\oplus W$, where $\dim(V)\geq T+1$ and $\dim(W)\geq m-1$. Any $x\in\mathbb{R}^{d}$ is written as $x=x_{V}+x_{W}$. Let $g:V\to\mathbb{R}$ be the hard scalar quadratic function from the previous proofs (either smooth convex or $\mu$-strongly convex as needed). In $W$, choose $m$ affinely independent points $\{a_{1},\dots,a_{m}\}$.
We define the objectives $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ for $i=1,\dots,m$ as:
$f_{i}(x)\;=\;g(x_{V})\;+\;\frac{\gamma}{2}\,\|x_{W}-a_{i}\|^{2}.$
The parameter $\gamma>0$ is chosen to ensure the desired smoothness and strong convexity.
  • Strongly convex case: Let $g$ be $L$-smooth and $\mu$-strongly convex. We set $\gamma=\mu$. The Hessian is $\nabla^{2}f_{i}(x)=\mathrm{diag}\bigl{(}\nabla^{2}g(x_{V}),\mu I_{W}\bigr{)}$. Since $\left\lVert\nabla^{2}g\right\rVert_{\mathrm{op}}\leq L$ and its eigenvalues are at least $\mu$, each $f_{i}$ is $L$-smooth and $\mu$-strongly convex.
  • Convex case: Let $g$ be $L$-smooth and convex. We set $\gamma=L$. The Hessian is $\nabla^{2}f_{i}(x)=\mathrm{diag}\bigl{(}\nabla^{2}g(x_{V}),LI_{W}\bigr{)}$. Since $\left\lVert\nabla^{2}g\right\rVert_{\mathrm{op}}\leq L$, each $f_{i}$ is $L$-smooth and convex.
The gradient is $\nabla f_{i}(x)=(\nabla g(x_{V}),\gamma(x_{W}-a_{i}))$. A convex combination of gradients is $\sum_{i}\lambda_{i}\nabla f_{i}(x)=\left(\nabla g(x_{V}),\gamma\left(x_{W}-\sum_{i}\lambda_{i}a_{i}\right)\right)$. The squared Pareto stationarity gap is therefore
$\mathcal{G}(x)^{2}\;=\;\min_{\lambda\in\Delta^{m}}\left\|\sum_{i}\lambda_{i}\nabla f_{i}(x)\right\|^{2}=\|\nabla g(x_{V})\|^{2}\;+\;\gamma^{2}\min_{\lambda\in\Delta^{m}}\left\|x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\right\|^{2}.$
This implies that $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$. A point $x$ is Pareto stationary ($\mathcal{G}(x)=0$) if and only if $\nabla g(x_{V})=0$ and $x_{W}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$. The Pareto set is thus $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{i}\}$, which is a non-singleton polytope.
We set the initial point to $x^{(0)}=x_{V}^{(0)}+a_{1}$, where $x_{V}^{(0)}$ is the starting point for the scalar hard instance, satisfying $\|x_{V}^{(0)}-x_{V}^{\star}\|=R$. The distance to the Pareto set is $\mathrm{dist}(x^{(0)},\mathcal{P})=\min_{p\in\mathcal{P}}\|x^{(0)}-p\|=\|(x_{V}^{(0)}+a_{1})-(x_{V}^{\star}+a_{1})\|=\|x_{V}^{(0)}-x_{V}^{\star}\|=R$.
Since the $V$-component of any gradient $\nabla f_{i}(x^{(j)})$ is always $\nabla g(x_{V}^{(j)})$, the update for $x_{V}$ is independent of $x_{W}$ and is identical to that of the algorithm run on the scalar problem $g$. Thus, the sequence $\{x_{V}^{(t)}\}$ is the hard sequence from the scalar lower bounds. Since $\mathcal{G}(x^{(t)})\geq\left\lVert\nabla g(x_{V}^{(t)})\right\rVert$, the scalar lower bounds apply directly to the stationarity gap of this non-degenerate MOO problem. ∎
5 Upper Bounds via Scalarization and Rate Comparison
The lower bounds can be matched by applying an optimal single-objective first-order method to a fixed scalarization of the multiobjective problem. Throughout this section, for any fixed $\lambda\in\Delta^{m}$ we consider the scalarized objective $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$, which is $L$-smooth (and $\mu$-strongly convex when stated). We write $x_{\lambda}^{\star}\in\arg\min f_{\lambda}$ and $R_{\lambda}:=\|x^{(0)}-x_{\lambda}^{\star}\|$.
To establish upper bounds that match the orders of our lower bounds, we now analyze the performance of an optimal first-order method applied to a fixed scalarization $f_{\lambda}$. We recall the celebrated accelerated gradient method (AGD) of Nesterov. The convergence guarantees for AGD are typically stated in terms of the objective function suboptimality, $f(x)-f^{*}$. Lemma 5.1 provides the crucial link needed to translate these guarantees into bounds on the gradient norm.
Lemma 5.1 (Descent lemma for the gradient norm).
If $f$ is an $L$-smooth convex function with minimum value $f^{\star}$, then for all $x$,
$\|\nabla f(x)\|^{2}\;\leq\;2L\big{(}f(x)-f^{\star}\big{)}.$
Definition 5.2 (AGD on a fixed scalarization).
Let $f_{\lambda}$ be $L$-smooth (and $\mu$-strongly convex when stated). We use the standard Nesterov accelerated gradient method:
  • Convex case ($\mu=0$). Initialize $x^{0}=y^{0}\in\mathbb{R}^{d}$ and $t_{0}=1$. For $k=0,1,\dots,T-1$,
    $x^{k+1}=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),\qquad t_{k+1}=\tfrac{1+\sqrt{1+4t_{k}^{\,2}}}{2},\qquad y^{k+1}=x^{k+1}+\tfrac{t_{k}-1}{t_{k+1}}\big{(}x^{k+1}-x^{k}\big{)}.$
  • Strongly convex case ($\mu>0$). Let $q:=\sqrt{\mu/L}$ and $\beta:=\tfrac{1-q}{1+q}$. Initialize $x^{0}=y^{0}\in\mathbb{R}^{d}$. For $k=0,1,\dots,T-1$,
    $x^{k+1}=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),\qquad y^{k+1}=x^{k+1}+\beta\big{(}x^{k+1}-x^{k}\big{)}.$
The convergence properties of this algorithm are well-established and provide optimal rates for the class of first-order methods on smooth convex and smooth, strongly convex functions. We summarize these classical results in the following lemma for completeness.
Lemma 5.3 (AGD convergence (see [24], §2.2; [5], §3.7.2; [4]).
Let $f$ be $L$-smooth and convex with minimizer $x^{\star}$ and $R:=\|x^{(0)}-x^{\star}\|$. The AGD iterates from Definition 5.2 satisfy
$f\big{(}x^{(T)}\big{)}-f(x^{\star})\;\leq\;\frac{2L\,R^{2}}{(T+1)^{2}}.$
If, in addition, $f$ is $\mu$-strongly convex with condition number $\kappa:=L/\mu$, then
$f\big{(}x^{(T)}\big{)}-f(x^{\star})\;\leq\;\frac{L+\mu}{2}\,R^{2}\,\Bigg{(}\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg{)}^{\!2T}.$
[…]
Theorem 5.4 (Upper bounds via AGD on a fixed scalarization).
Assume each $f_{i}$ is convex and $L$-smooth. Fix any $\lambda\in\Delta^{m}$ and run AGD (Definition 5.2) on $f_{\lambda}$ for $T$ steps from $x^{(0)}$. Then
$\mathcal{G}\big{(}x^{(T)}\big{)}\;\leq\;\|\nabla f_{\lambda}(x^{(T)})\|\;\leq\;\frac{2LR_{\lambda}}{T+1}.$
If, moreover, each $f_{i}$ is $\mu$-strongly convex (so $f_{\lambda}$ is $\mu$-strongly convex with $\kappa=L/\mu$), then
$\mathcal{G}\big{(}x^{(T)}\big{)}\;\leq\;\|\nabla f_{\lambda}(x^{(T)})\|\;\leq\;\sqrt{\,L(L+\mu)\,}\;R_{\lambda}\;\Bigg{(}\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg{)}^{\!T}.$
Proof.
For any $x$, $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$ by definition of the Pareto stationarity gap. In the convex case, combine Lemma 5.1 with Lemma 5.3 applied to $f_{\lambda}$:
$\|\nabla f_{\lambda}(x^{(T)})\|^{2}\;\leq\;2L\big{(}f_{\lambda}(x^{(T)})-f_{\lambda}(x_{\lambda}^{\star})\big{)}\;\leq\;2L\cdot\frac{2LR_{\lambda}^{2}}{(T+1)^{2}}\;=\;\frac{4L^{2}R_{\lambda}^{2}}{(T+1)^{2}},$
and take square roots. In the strongly convex case, apply the same inequality $\|\nabla f_{\lambda}(x)\|^{2}\leq 2L(f_{\lambda}(x)-f_{\lambda}^{\star})$ together with the strongly convex part of Lemma 5.3:
$\|\nabla f_{\lambda}(x^{(T)})\|^{2}\;\leq\;2L\cdot\frac{L+\mu}{2}\,R_{\lambda}^{2}\,\Bigg{(}\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg{)}^{\!2T}\;=\;L(L+\mu)\,R_{\lambda}^{2}\,\Bigg{(}\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg{)}^{\!2T},$
and take square roots to obtain the stated bound. Note that AGD is an oblivious span method (but not an oblivious one-step method). Hence an $\mathcal{O}(1/T)$ upper bound in the convex case is achievable by an oblivious span method on a fixed scalarization; the linear bound in the strongly convex case follows analogously with the classical accelerated factor. ∎
Corollary 5.5 (AGD $\varepsilon$-complexity for the Pareto gap).
Under the assumptions of Theorem 5.4 (convex case), to guarantee $\mathcal{G}(x^{(T)})\leq\varepsilon$ it suffices to take
[…]
In the $\mu$-strongly convex case with $\kappa=L/\mu$, it suffices to take
[…]
that is, $T=\Theta\!\big{(}\sqrt{\kappa}\,\log(\sqrt{L(L+\mu)}\,R_{\lambda}/\varepsilon)\big{)}$.
Remark 5.6 (Best known one-step upper bound).
For $L$-smooth convex $f_{\lambda}$ and stepsizes $\alpha_{t}\leq 1/L$, gradient descent satisfies
$\sum_{t=0}^{T-1}\|\nabla f_{\lambda}(x^{(t)})\|^{2}\;\leq\;2L\big{(}f_{\lambda}(x^{(0)})-f_{\lambda}^{\star}\big{)}\;\leq\;L^{2}R_{\lambda}^{2},$
whence $\min_{0\leq t<T}\|\nabla f_{\lambda}(x^{(t)})\|\leq LR_{\lambda}/\sqrt{T}$. Since $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$, we obtain $\min_{0\leq t<T}\mathcal{G}(x^{(t)})\leq LR_{\lambda}/\sqrt{T}$ for the oblivious one-step class; this is tight in general for that class.
Remark 5.7 (Instance calibration and rate comparison).
The upper bounds above depend on $R_{\lambda}=\|x^{(0)}-x_{\lambda}^{\star}\|$, whereas our lower bounds depend on $R=\mathrm{dist}(x^{(0)},\mathcal{P})$. Always $R\leq R_{\lambda}$. For the hard instances in Theorem 4.5, choosing $\lambda=e_{1}$ yields $R_{e_{1}}=R$, so the orders match. Importantly, the $\Omega(1/T)$ lower bound in Theorem 4.3 holds for the oblivious one-step class, whereas the $\mathcal{O}(1/T)$ upper bound here is achieved by an oblivious span method (AGD). Whether $\mathcal{O}(1/T)$ is achievable within the oblivious one-step class remains open; the best known upper bound there is $\mathcal{O}(LR/\sqrt{T})$.
6 Conclusion
We developed an oracle complexity framework for smooth multiobjective optimization (MOO) focused on the task of finding $\varepsilon$-Pareto stationary points. Our lower bounds cover both convex and strongly convex regimes and distinguish between algorithm classes. In the $\mu$-strongly convex case, we proved a tight linear-rate lower bound that implies a worst-case iteration complexity of $\Theta\!\big{(}\sqrt{\kappa}\log(1/\varepsilon)\big{)}$ with $\kappa=L/\mu$, matching classical single-objective theory when specialized to scalarizations. In the merely convex case, we established an $\Omega(1/T)$ min-iterate lower bound for oblivious one-step methods and a universal $\Omega\!\big{(}1/T^{2}\big{)}$ last-iterate lower bound for fully adaptive span first-order methods, thereby clarifying the landscape across non-adaptive and adaptive algorithms.
On the positive side, we showed that applying an optimal scalar method (e.g., AGD) to a fixed scalarization attains an $O(1/T)$ upper bound for the Pareto stationarity gap; this matches our $\Omega(1/T)$ rate up to the difference in algorithm classes (oblivious span vs. oblivious one-step). We also calibrated upper and lower bounds via the initialization radii $R$ and $R_{\lambda}$ and proved that our constructions are robust: they apply to instances with genuinely distinct objectives and non-singleton Pareto fronts. Altogether, the results position MOO on a footing analogous to single-objective optimization while revealing new, vector-specific phenomena.
Two main directions remain open. First, in the convex setting for adaptive span methods, there is a gap between our universal $\Omega\!\big{(}1/T^{2}\big{)}$ last-iterate lower bound and known $O(1/T)$ upper bounds: closing this — either by stronger lower bounds (e.g., universal $\Omega(1/T)$) or by matching algorithms under the same oracle model — is a central challenge. Second, extending the framework to broader models (stochastic oracles, inexact gradients, constraints/composite objectives, and nonconvex MOO), as well as investigating higher-order information and dimension dependence, would further refine the fundamental limits that govern multiobjective first-order methods.
Appendix A Appendix: Derivations of Polynomial Extremal Results
A.1 A tight $\Omega(1/T)$ lower bound for oblivious one-step methods (with $L$-capped steps)
In this appendix we establish a tight $\Omega(1/T)$ lower bound for the stationarity-gap proxy
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)},$
which arises in the analysis of oblivious one-step methods on $L$-smooth convex quadratics. The proof is self-contained and uses only elementary inequalities.
Model and class.
We consider oblivious one-step gradient methods applied to a fixed scalarization $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ of a multiobjective instance, where $H\succeq 0$ and $\sigma(H)\subset[0,L]$. The method generates
$x^{(t+1)}\;=\;x^{(t)}-\alpha_{t}\,\nabla g\bigl{(}x^{(t)}\bigr{)},\qquad t=0,1,2,\dots,$
with a pre-scheduled stepsize sequence $\{\alpha_{t}\}_{t\geq 0}$ which may depend on known problem data (here, $L$) but not on oracle feedback. We adopt the natural $L$-cap:
$0\;\leq\;\alpha_{t}\;\leq\;\frac{1}{L}\qquad\text{for all }t\geq 0.$ (9)
This is the standard safe region for gradient descent on $L$-smooth convex quadratics with known $L$ and is the relevant oblivious regime for worst-case analysis.
Polynomial representation.
Let $e^{(t)}=x^{(t)}-x^{\star}$, where $x^{\star}\in\arg\min g$; then for quadratics $e^{(t)}=p_{t}(H)\,e^{(0)}$ with
$p_{t}(\zeta)\;=\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)},\qquad p_{0}\equiv 1,$
and hence $\nabla g\bigl{(}x^{(t)}\bigr{)}=H\,p_{t}(H)\,e^{(0)}$ (Lemma 3.13).
We prove the following extremal result.
Theorem A.1 (Tight $\Omega(1/T)$ product-form extremal under $L$-capped steps).
Fix $L>0$ and $t\in\mathbb{N}$. For any nonnegative stepsizes $\alpha_{0},\ldots,\alpha_{t-1}$ satisfying (9), we have
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)}\;\;\geq\;\;\frac{L}{4\,(t+1)}.$ (10)
Moreover, the bound is tight up to a universal constant: the constant stepsizes $\alpha_{k}\equiv 1/L$ yield
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)}\;=\;\max_{\zeta\in[0,L]}\;\zeta\Bigl{(}1-\frac{\zeta}{L}\Bigr{)}^{t}\;\;\leq\;\;\frac{L}{e\,(t+1)}.$ (11)
Remark A.2 (Scope of (9)).
The cap (9) is standard in first-order analysis with known $L$. Allowing stepsizes larger than $1/L$ may cause ascent on $L$-smooth convex quadratics and does not help the algorithm in the worst case. Our lower bound therefore covers the canonical oblivious regime.
The proof rests on a simple but sharp product inequality.
Lemma A.3 (A product lower bound).
For any $r\in\mathbb{N}$ and any numbers $x_{1},\ldots,x_{r}\in[0,1]$,
$\prod_{i=1}^{r}\bigl{(}1-x_{i}\bigr{)}\;\;\geq\;\;1-\sum_{i=1}^{r}x_{i}.$ (12)
Proof.
By induction on $r$. The case $r=1$ is trivial. Suppose (12) holds for r1r-1. Then for $r$ we have
$\prod_{i=1}^{r}(1-x_{i})\;=\;\Bigl{(}\prod_{i=1}^{r-1}(1-x_{i})\Bigr{)}(1-x_{r})\;\geq\;\Bigl{(}1-\sum_{i=1}^{r-1}x_{i}\Bigr{)}(1-x_{r})\;=\;1-\sum_{i=1}^{r}x_{i}\;+\;x_{r}\sum_{i=1}^{r-1}x_{i}\;\geq\;1-\sum_{i=1}^{r}x_{i},$
since $x_{r}\sum_{i=1}^{r-1}x_{i}\geq 0$. ∎
Proof of Theorem A.1.
Fix $t\geq 1$ and a stepsize sequence $\alpha_{0},\ldots,\alpha_{t-1}$ with $0\leq\alpha_{k}\leq 1/L$ for all $k$. For any $\zeta\in[0,L]$, set $x_{k}:=\alpha_{k}\,\zeta\in[0,1]$. Applying Lemma A.3 gives
$\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)}\;\;\geq\;\;1-\zeta\sum_{k=0}^{t-1}\alpha_{k}.$ (13)
Define the partial sums $S_{t}:=\sum_{k=0}^{t-1}\alpha_{k}$. Note that $S_{t}\leq t\cdot(1/L)=t/L$ by (9).
Consider the function $\Phi_{t}(\zeta):=\zeta\prod_{k=0}^{t-1}(1-\alpha_{k}\,\zeta)$ on $[0,L]$. By (13),
$\Phi_{t}(\zeta)\;\geq\;\zeta\bigl{(}1-\zeta S_{t}\bigr{)}\qquad\text{for all }\zeta\in[0,L].$ (14)
Let $\zeta^{\star}:=\min\{L,\;1/(2S_{t})\}$ (with the convention $1/0=+\infty$). We distinguish two exhaustive cases.
Case 1: $S_{t}\geq 1/(2L)$. Then $\zeta^{\star}=1/(2S_{t})\leq L$, and hence, by (14),
$\Phi_{t}(\zeta^{\star})\;\geq\;\zeta^{\star}\Bigl{(}1-\zeta^{\star}S_{t}\Bigr{)}\;=\;\frac{1}{2S_{t}}\cdot\Bigl{(}1-\tfrac{1}{2}\Bigr{)}\;=\;\frac{1}{4S_{t}}\;\geq\;\frac{1}{4\,(t/L)}\;=\;\frac{L}{4t}.$
Case 2: $S_{t}<1/(2L)$. Then $\zeta^{\star}=L$ and, by (14),
$\Phi_{t}(\zeta^{\star})\;\geq\;L\bigl{(}1-LS_{t}\bigr{)}\;\geq\;L\Bigl{(}1-\tfrac{1}{2}\Bigr{)}\;=\;\frac{L}{2}\;\geq\;\frac{L}{4t}\qquad(\text{since }t\geq 1).$
In both cases, we obtain $\max_{\zeta\in[0,L]}\Phi_{t}(\zeta)\geq L/(4t)$. Since $t\geq 1$ implies $1/t\geq 1/(t+1)$, this establishes the slightly looser bound in (10).
For the near-matching upper bound (11), take the constant stepsizes $\alpha_{k}\equiv 1/L$. Then
$\Psi_{t}(\zeta):=\zeta\Bigl{(}1-\frac{\zeta}{L}\Bigr{)}^{t},\qquad\zeta\in[0,L].$
Elementary calculus shows that $\Psi_{t}$ attains its maximum at $\zeta^{\diamond}=L/(t+1)$, with
$\max_{\zeta\in[0,L]}\Psi_{t}(\zeta)\;=\;\Psi_{t}(\zeta^{\diamond})\;=\;\frac{L}{t+1}\left(1-\frac{1}{t+1}\right)^{t}\;\leq\;\frac{L}{t+1}\,\mathrm{e}^{-\,t/(t+1)}\;\leq\;\frac{L}{\mathrm{e}\,(t+1)}.$
This establishes (11). ∎
We now translate Theorem A.1 into an oracle lower bound for the gradient norm.
Corollary A.4 (Oblivious one-step lower bound for the gradient norm).
Let $t\in\mathbb{N}$ and let an oblivious one-step method with stepsizes $\alpha_{0},\ldots,\alpha_{t-1}$ obeying (9) be given. Then there exists an $L$-smooth convex quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ with $\sigma(H)\subset[0,L]$ and an initialization $x^{(0)}$ satisfying $\|x^{(0)}-x^{\star}\|=R$ such that
$\min_{0\leq s\leq t}\;\bigl{\|}\nabla g\bigl{(}x^{(s)}\bigr{)}\bigr{\|}\;\;\geq\;\;\frac{L\,R}{4\,(t+1)}.$
Proof.
For quadratics, $\|\nabla g(x^{(s)})\|=\|H\,p_{s}(H)\,(x^{(0)}-x^{\star})\|$. Fix any unit vector $u$ and take $x^{(0)}-x^{\star}=R\,u$. For every fixed $\zeta\in[0,L]$, the monotonicity $(1-\alpha_{s}\zeta)\in[0,1]$ implies
$\zeta\,\Bigl{|}\prod_{k=0}^{s-1}\bigl{(}1-\alpha_{k}\zeta\bigr{)}\Bigr{|}\;\;\text{is nonincreasing in $s$.}$
Hence, for any fixed $\zeta$,
$\min_{0\leq s\leq t}\ \zeta\,\Bigl{|}\prod_{k=0}^{s-1}\bigl{(}1-\alpha_{k}\zeta\bigr{)}\Bigr{|}\;=\;\zeta\,\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\zeta\bigr{)}.$
Choosing $H$ to be diagonal with an eigenvalue at the maximizer $\zeta^{\star}\in[0,L]$ of the right-hand side and aligning $u$ with the corresponding eigenvector yields
$\min_{0\leq s\leq t}\;\|\nabla g(x^{(s)})\|\;=\;R\cdot\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl{(}1-\alpha_{k}\,\zeta\bigr{)}\;\geq\;\frac{L\,R}{4\,(t+1)},$
by Theorem A.1. ∎
This result is lifted to the multiobjective setting using the construction from Theorem 4.5. For the instance $f_{i}(x)=g(x_{V})+\frac{L}{2}\|x_{W}-a_{i}\|^{2}$, we have $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$. Corollary A.4 thus directly implies a lower bound on $\min_{s\leq t}\mathcal{G}(x^{(s)})$. This establishes the bound for a non-degenerate MOO instance with distinct objectives and a non-singleton Pareto set.
A.2 Universal Markov-Based Bound for Span Methods
The lower bound for adaptive methods relies on $\inf_{p_{t}}\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|$ where $p_{t}$ is any polynomial with $\deg p_{t}\leq t$ and $p_{t}(0)=1$.
Lemma A.5 (Markov Brothers’ Inequality (see [28], Chapter 2, Section 2.7, Page 123)).
For any polynomial $q(x)$ of degree $n$, $\max_{x\in[-1,1]}|q^{\prime}(x)|\leq n^{2}\max_{x\in[-1,1]}|q(x)|$.
Let $S(\zeta)=\zeta p_{t}(\zeta)$, which is a polynomial of degree $t+1$. We know $S(0)=0$ and from $p_{t}(0)=1$, we have $S^{\prime}(0)=p_{t}(0)+0\cdot p_{t}^{\prime}(0)=1$. Let’s map the interval $[0,L]$ to [1,1][-1,1] via the transformation $x=\frac{2\zeta}{L}-1$, which means $\zeta=\frac{L(x+1)}{2}$. Define a new polynomial $q(x)=S\left(\frac{L(x+1)}{2}\right)$. The degree of $q$ is $n=t+1$. At x=1x=-1, q(1)=S(0)=0q(-1)=S(0)=0. The derivative of $q$ with respect to $x$ is $q^{\prime}(x)=S^{\prime}\left(\frac{L(x+1)}{2}\right)\cdot\frac{L}{2}$. At x=1x=-1, this gives $q^{\prime}(-1)=S^{\prime}(0)\cdot\frac{L}{2}=1\cdot\frac{L}{2}=\frac{L}{2}$. By Markov’s inequality:
$\frac{L}{2}=|q^{\prime}(-1)|\leq\max_{x\in[-1,1]}|q^{\prime}(x)|\leq(t+1)^{2}\max_{x\in[-1,1]}|q(x)|.$
Rearranging gives the bound on the maximum of $q$, which is the same as the maximum of $S$:
$\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|=\max_{x\in[-1,1]}|q(x)|\geq\frac{L}{2(t+1)^{2}}.$
A.3 Chebyshev Extremal Problem on $\mu,L]$
The lower bound for strongly convex methods requires solving $\min_{p_{t}}\max_{\zeta\in[\mu,L]}|p_{t}(\zeta)|$ subject to $\deg p_{t}\leq t$ and $p_{t}(0)=1$. The affine transformation $\xi=\frac{2\zeta-(L+\mu)}{L-\mu}$ maps the interval $\zeta\in[\mu,L]$ to $\xi\in[-1,1]$. The constraint $p_{t}(0)=1$ is evaluated at $\zeta=0$, which corresponds to $\xi_{0}=-\frac{L+\mu}{L-\mu}$. Let $\tilde{p}_{t}(\xi)=p_{t}(\zeta(\xi))$. The problem becomes:
$\min_{\begin{subarray}{c}\deg\tilde{p}_{t}\leq t\\ \tilde{p}_{t}(\xi_{0})=1\end{subarray}}\max_{\xi\in[-1,1]}|\tilde{p}_{t}(\xi)|.$
The solution to this classic problem is the scaled Chebyshev polynomial $\tilde{p}_{t}^{*}(\xi)=\frac{T_{t}(\xi)}{T_{t}(\xi_{0})}$, where $T_{t}$ is the Chebyshev polynomial of the first kind of degree $t$ (see [28], Chapter 2, Section 2.7, Page 108). The extremal value is $\frac{1}{|T_{t}(\xi_{0})|}$ because $\max_{\xi\in[-1,1]}|T_{t}(\xi)|=1$. With $\kappa=L/\mu$, we have $|\xi_{0}|=\frac{L+\mu}{L-\mu}=\frac{\kappa+1}{\kappa-1}>1$. For $|x|>1$, the Chebyshev polynomial is given by $T_{t}(x)=\frac{1}{2}\left((x+\sqrt{x^{2}-1})^{t}+(x-\sqrt{x^{2}-1})^{-t}\right)$. Let $\rho=|\xi_{0}|+\sqrt{|\xi_{0}|^{2}-1}$. A direct calculation shows:
$\rho=\frac{\kappa+1}{\kappa-1}+\sqrt{\left(\frac{\kappa+1}{\kappa-1}\right)^{2}-1}=\frac{\kappa+1+2\sqrt{\kappa}}{\kappa-1}=\frac{(\sqrt{\kappa}+1)^{2}}{(\sqrt{\kappa}-1)(\sqrt{\kappa}+1)}=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
Then $|T_{t}(\xi_{0})|=|T_{t}(|\xi_{0}|)|=\frac{1}{2}(\rho^{t}+\rho^{-t})$. The extremal value is $\frac{1}{|T_{t}(\xi_{0})|}=\frac{2}{\rho^{t}+\rho^{-t}}$.