### v2

# Complexity Bounds for Smooth Multiobjective Optimization

Abstract
We study the oracle complexity of finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization with $m$ objectives. Progress is measured by the Pareto stationarity gap $\mathcal{G}(x)$, the norm of the best convex combination of objective gradients. Our analysis relies on a non-degenerate lifting that embeds hard single-objective instances into MOO instances with distinct objectives and non-singleton Pareto fronts while preserving lower bounds on $\mathcal{G}$. We establish: (i) in the $\mu$-strongly convex case, any span first-order method has worst-case linear convergence no faster than $\exp(-\Theta(T/\sqrt{\kappa}))$ after $T$ oracle calls, yielding $\Theta(\sqrt{\kappa}\log(1/\varepsilon))$ iterations and matching accelerated upper bounds; (ii) in the convex case, an $\Omega(1/T)$ min-iterate lower bound for oblivious one-step methods and a universal last-iterate lower bound $\Omega(1/T^{2})$ for oblivious span methods via polynomial-degree arguments, and we further show this latter bound is loose (for general adaptive methods) by importing geometric lower bounds to obtain an $\Omega(1/T)$ min-iterate lower bound for general adaptive first-order methods; (iii) in the nonconvex case with $L$-Lipschitz gradients, an $\Omega(\sqrt{L}/(T+1))$-type lower bound on $\mathcal{G}$ (tight in order), implying $\Omega(1/\varepsilon^{2})$ iterations to reach $\mathcal{G}(x)\leq\varepsilon$ up to natural scaling.
Keywords Multiobjective optimization; Oracle complexity; Pareto stationarity; First-order span methods; Adaptive scalarization
1 Introduction
Multiobjective optimization (MOO) provides a mathematical framework for decision-making problems involving multiple, often conflicting, performance criteria. Given a vector-valued objective function $F(x)=(f_{1}(x),\dots,f_{m}(x))$, the goal is to find a point $x$ that offers the best possible trade-off among the individual objectives $f_{i}$. Since a single point that simultaneously minimizes all objectives typically does not exist, the central solution concept is that of Pareto optimality. A point is Pareto optimal if no other point can improve one objective without degrading at least one other.
First-order iterative methods have become the tool of choice for solving large-scale MOO problems, particularly in machine learning [33], and they are increasingly deployed in engineering and finance — for example, distributed-gradient schemes for multi-objective AC optimal power flow [26], adjoint-based gradient methods for reservoir well-control under risk and uncertainty [23], and gradient-descent formulations for multi-objective portfolio selection [30]. These methods rely on an oracle that returns function values and gradients. A key challenge is to define a tractable computational goal. While computing an exact Pareto-optimal point is often impractical, a widely adopted necessary condition is Pareto criticality [15], which is equivalent to the Pareto stationarity notion adopted here. A point $x$ is Pareto stationary if the convex hull of its gradients contains the origin. The proximity to this condition is measured by the Pareto stationarity gap $\mathcal{G}(x)$ (see Definition 3.5 in Section 3). Finding a point $x$ with $\mathcal{G}(x)\leq\varepsilon$ is a standard objective for many contemporary MOO algorithms [12].
Although numerous algorithms have been proposed with convergence-rate guarantees (upper bounds), the fundamental limits of what is achievable have remained largely unexplored. Without oracle lower bounds, we cannot know if existing algorithms are optimal. In single-objective optimization, this story is complete: the seminal work of Nemirovski and Yudin [27] established lower bounds that were later matched by accelerated methods [28], proving their optimality. For MOO, however, a corresponding theory of oracle complexity has been absent.
Why scalarization does not trivialize MOO
It may appear that multiobjective complexity reduces immediately to the single-objective case by fixing a scalarization $f_{\lambda}=\sum_{i}\lambda_{i}f_{i}$ and applying known convergence results. Indeed, all classical rates for smooth convex or strongly convex minimization (e.g., gradient descent $O(LR/\sqrt{T})$, AGD $O(LR/T)$, and linear convergence under strong convexity) apply verbatim to any fixed scalarization. However, the multiobjective goal is different: the central complexity measure is the Pareto stationarity gap $\mathcal{G}(x)$ which requires considering all convex combinations of gradients simultaneously. A point minimizing a fixed $f_{\lambda}$ need not minimize $\mathcal{G}$, and the algorithm classes we study (oblivious one-step vs. oblivious span vs. general adaptive first-order methods) differ in their ability to exploit information across objectives. Consequently, while scalarization provides useful upper bounds, the oracle complexity of Pareto stationarity requires a dedicated analysis, which is the focus of this paper.
Contributions
This paper closes this fundamental gap by providing a nuanced information-based complexity analysis that distinguishes between different classes of first-order algorithms. Our main contributions are as follows:
  • Robust non-degenerate lifting. We show our lower bounds are not artifacts of degenerate problems. We construct MOO instances with distinct objectives and non-singleton Pareto fronts that preserve lower bounds on $\mathcal{G}$ from embedded scalar instances.
  • Tight bounds for strongly convex MOO: We establish a tight linear convergence lower bound. Any span first-order method can be forced to have a stationarity gap of at least $\mu R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T}$, implying an iteration complexity of $\Omega(\sqrt{\kappa}\log(\mu R/\varepsilon))$. This is matched by accelerated gradient methods on a fixed scalarization.
  • Lower bound for oblivious one-step convex MOO: For the class of oblivious one-step gradient methods—which includes gradient descent with a pre-scheduled step-size sequence—we prove an $\Omega(LR/(T+1))$ lower bound on the (min-iterate) stationarity gap. Accelerated gradient descent (AGD), which is not in this class but is an oblivious span method, achieves an $\mathcal{O}(LR/(T+1))$ upper rate on a fixed scalarization. Thus the $\Omega(1/T)$ rate is matched by oblivious span methods, while within the oblivious one-step class the best known general upper bound is $\mathcal{O}(LR/\sqrt{T})$.
  • Polynomial-degree lower bound for oblivious span methods and geometric lower bound for adaptive methods: Within the oblivious span model, we establish a universal last-iterate polynomial-degree lower bound $\Omega(1/T^{2})$ via Markov-type extremal inequalities. For general deterministic adaptive first-order methods, we prove a stronger geometric (zero-chain) $\Omega(1/T)$ min-iterate lower bound under a bounded scalarization gap, obtained by lifting the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [7].
  • Nonconvex extension with Lipschitz gradients. Using the same lifting on the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [6], we obtain a nonconvex MOO lower bound under $L$-Lipschitz gradients of $\Omega(\sqrt{L\Delta/(T+1)})$ (equivalently $\Omega(L\Delta/\varepsilon^{2})$ iterations) to reach $\mathcal{G}(x)\leq\varepsilon$.
Our methodology combines classical polynomial-based lower-bound techniques in the spirit of the resisting-oracle framework [27] with a non-degenerate lifting reduction that transfers sharp single-objective stationary-point lower bounds (e.g., [7]) to multiobjective Pareto stationarity. This yields a unified oracle-complexity perspective for smooth MOO and clarifies how algorithmic restrictions (oblivious one-step vs. oblivious span vs. general adaptive first-order methods) and iterate selection (min vs. last) shape the achievable rates.
Organization
Section 2 reviews related work. Section 3 fixes notation, defines $\mathcal{G}$, and formalizes oracle models and algorithm classes. Section 4 presents the lifting construction, then derives strongly convex, convex, and nonconvex lower bounds by combining polynomial extremal arguments and geometric stationary-point lower bounds. Section 5 provides matching upper bounds via AGD on fixed scalarizations and compares rates against the lower bounds. Section 6 concludes with implications and open questions. The appendices contain the polynomial extremal derivations (Chebyshev and Markov) used to analyze span and oblivious methods on hard quadratic instances.
2 Related Work
Oracle complexity and first-order limits
The information-based complexity program for single-objective convex optimization established sharp first-order lower bounds together with matching optimal methods [27, 28]. In [27], these lower bounds are derived by reducing the worst-case error to an extremal polynomial problem, solved exactly by (scaled) Chebyshev polynomials; see [32] for the underlying Chebyshev minimax and related inequalities. Recent analyses include Performance Estimation Problem (PEP) frameworks that deliver tight worst-case guarantees for many first-order methods [10]. Beyond the classical oracle model, a line of work formalizes oblivious and structured first-order algorithms: Arjevani and Shamir [3] introduced the oblivious framework with dimension-free lower bounds for $L$-smooth convex and $L$-smooth $\mu$-strongly convex problems; Arjevani, Shalev-Shwartz, and Shamir [1] developed the $p$-SCLI framework (covering p=1p{=}1 one-step methods) with matching bounds; and Arjevani and Shamir [2] extended dimension-free iteration lower bounds to finite-sum settings, covering variance-reduced families. Complementary results for randomized first-order algorithms appear in Woodworth and Srebro [36]. More recently, Carmon, Duchi, Hinder, and Sidford [7] established sharp lower bounds for finding first-order stationary points in smooth convex and nonconvex optimization, clarifying the fundamental $1/T$ (convex) and $1/\sqrt{T}$ (nonconvex) barriers in gradient-norm complexity under standard first-order oracles. A comprehensive and recent treatment of finite-sum lower bounds is given by Han, Xie, and Zhang [21].
Multiobjective optimization algorithms
Classical foundations and broad context are covered in the monograph by Miettinen [25] and the survey of Marler and Arora [24]. For descent-type methods, the steepest–descent family beginning with [15] and including cone-ordered variants and refined convergence analyses [11, 9] provides global convergence to Pareto-stationary points. Worst-case complexity guarantees of order $\mathcal{O}(\varepsilon^{-2})$ for trust-region and regularization frameworks in unconstrained MOO were established in [20]. Complexity aspects for gradient descent in smooth MOO are analyzed in [13]. Universal nonmonotone line-search frameworks for non-convex MOO with convex constraints have also been proposed [31]. When progress is tracked through standard merit functions or stationarity surrogates, the resulting upper bounds are typically sublinear (e.g., $\mathcal{O}(1/k)$) in smooth convex settings [35, 18]. Second-order frameworks — Newton-type and SQP — offer globalization with fast local behavior: globally convergent Newton methods for MOO with fast (often superlinear) local rates under standard regularity are developed in [19], while SQP-based schemes for constrained problems are given in [14]. Conditional-gradient (Frank–Wolfe) methods have been developed for vector/multiobjective optimization with global convergence to weakly efficient (Pareto-stationary) solutions under cone-convexity assumptions [8]. Away-step and related variants can accelerate convergence and, under additional geometric or curvature assumptions, deliver faster (sometimes linear) rates [17, 18], with adaptive generalized conditional-gradient methods extending this line [16]. Accelerated first-order schemes tailored to the multiobjective setting continue to mature: an accelerated proximal-gradient method with $\mathcal{O}(1/k^{2})$ decrease for an appropriate merit measure is established in [34], and recent work clarifies monotonicity properties of multiobjective APG [29]. Complementary practical gradient-based schemes and stationarity criteria, such as multiple-gradient descent, are surveyed in [12], and links to multi-task learning further motivate stationarity-based goals and algorithm design [33].
Our position
While scalarization-based upper bounds are classical, a systematic oracle-complexity theory for Pareto stationarity that (i) distinguishes oblivious one-step from both oblivious span methods and general adaptive first-order methods, (ii) provides robust non-degenerate hard MOO instances, and (iii) connects polynomial-degree lower bounds to geometric stationary-point lower bounds, has been missing. We supply such a framework and import sharp scalar stationary-point lower bounds into MOO via an explicit lifting reduction.
3 Preliminaries
3.1 Notation
All vectors lie in $\mathbb{R}^{d}$ and matrices in $\mathbb{R}^{d\times d}$. The Euclidean inner product and norm are $\langle x,y\rangle$ and $\|x\|$; for a matrix $A$, $\|A\|_{\mathrm{op}}$ is the operator norm, $\sigma(A)$ its spectrum, $A\succeq 0$ (resp. $A\succ 0$) means positive (semi)definite, and $I$ is the identity; $(\cdot)^{\top}$ denotes transpose. For a set $S$, $\operatorname{conv}(S)$ is its convex hull, $\operatorname{span}(S)$ its linear span, and $\operatorname{dist}(x,S):=\inf_{y\in S}\|x-y\|$ its distance. For a matrix $A$, $\operatorname{range}(A)$ and $\ker(A)$ denote its range and kernel. We write $\mathbb{N}_{0}:=\{0,1,2,\ldots\}$ and $\mathbb{N}:=\{1,2,\ldots\}$.
For $m\geq 2$ objectives $F(x)=(f_{1}(x),\dots,f_{m}(x))$, the unit simplex is
$\Delta^{m}:=\{\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\}.$
For $\lambda\in\Delta^{m}$, the scalarization is $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ and $x^{\star}_{\lambda}\in\arg\min_{x}f_{\lambda}(x)$ denotes an arbitrary minimizer.
In convex settings we use the scalarization value gap at initialization
$\Delta_{\lambda}(x^{(0)}):=f_{\lambda}(x^{(0)})-\inf_{x}f_{\lambda}(x).$
When the Pareto set $\mathcal{P}$ is well-defined (convex case; see Proposition 3.8), we write
$R:=\mathrm{dist}(x^{(0)},\mathcal{P}).$
In the nonconvex extension we will instead refer to the set of Pareto-stationary points $\mathcal{S}:=\{x:\mathcal{G}(x)=0\}$ when needed.
Iterates are $x^{(t)}$ for $t\in\mathbb{N}_{0}$. $T\in\mathbb{N}$ is the number of oracle calls/iterations.
The condition number is $\kappa:=L/\mu$ when strong convexity applies. We occasionally write $\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}$ in spectral formulas.
For quadratics $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, we use the spectral variable $\zeta$ and real polynomials $p_{t}$ (with $\deg p_{t}\leq t$, $p_{t}(0)=1$) to express errors via $x^{(t)}-x^{\star}=p_{t}(H)\bigl(x^{(0)}-x^{\star}\bigr)$. Chebyshev polynomials of the first kind are denoted $T_{t}(\cdot)$.
Standard Landau symbols $O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ are used; $\log$ denotes the natural logarithm.
For nonnegative quantities $A$ and $B$, we write $A\asymp B$ (equivalence up to universal constants) if there exist absolute constants $c_{1},c_{2}>0$ (independent of problem parameters such as $L,\Delta,R,T,m,d$) such that $c_{1}B\leq A\leq c_{2}B$. We use this notation only when the implicit constants are universal; otherwise we state explicitly which parameters they may depend on.
The vectors $(e_{i})_{i=1}^{m}$ are the canonical basis of $\mathbb{R}^{m}$.
We consider unconstrained optimization problems in $\mathbb{R}^{d}$. A differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth if its gradient is $L$-Lipschitz continuous. It is $\mu$-strongly convex if $f(y)\geq f(x)+\left\langle\nabla f(x),y-x\right\rangle+\frac{\mu}{2}\left\lVert y-x\right\rVert^{2}$ for all $x,y$.
3.2 Problem formulation
We consider unconstrained multiobjective problems
$\min_{x\in\mathbb{R}^{d}}F(x):=(f_{1}(x),\ldots,f_{m}(x)).$ (1)
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable. Convexity and strong convexity will be imposed in the relevant sections.
3.3 Pareto optimality and first-order conditions
We now review the standard optimality notions for (possibly non-convex) MOO problems (see [25]) and clarify their relationship with the Pareto stationarity gap used throughout the paper.
Definition 3.1 (Dominance).
Given $x,y\in\mathbb{R}^{d}$, we say that $y$ dominates $x$ if
$f_{i}(y)\leq f_{i}(x)\ \ \text{for all }i=1,\dots,m,\quad\text{and}\quad f_{j}(y)<f_{j}(x)\ \ \text{for at least one }j.$
Based on this, a point is considered optimal if no other point in the domain dominates it.
Definition 3.2 (Pareto optimality (strong)).
A point $x^{\star}\in\mathbb{R}^{d}$ is Pareto optimal if there is no $y\in\mathbb{R}^{d}$ that dominates $x^{\star}$; equivalently, there is no $y$ with
$f_{i}(y)\leq f_{i}(x^{\star})\ \ \text{for all }i,\quad\text{and}\quad f_{j}(y)<f_{j}(x^{\star})\ \ \text{for some }j.$
A related, slightly weaker condition is often useful for theoretical analysis and algorithm design:
Definition 3.3 (Weak Pareto optimality).
A point $x^{\star}\in\mathbb{R}^{d}$ is weakly Pareto optimal if there is no $y\in\mathbb{R}^{d}$ such that
$f_{i}(y)<f_{i}(x^{\star})\ \ \text{for all }i=1,\dots,m.$
To develop iterative, gradient-based methods, we need a computable, first-order necessary condition for optimality. This leads to the concept of Pareto criticality.
Definition 3.4 (Pareto criticality).
Following [15], a point $x\in\mathbb{R}^{d}$ is called Pareto critical if there is no common strict descent direction, i.e.,
$\not\exists\,v\in\mathbb{R}^{d}\quad\text{such that}\quad\nabla f_{i}(x)^{\top}v<0\quad\text{for all }i=1,\dots,m.$
Equivalently, with the Jacobian $JF(x)\in\mathbb{R}^{m\times d}$ whose $i$-th row is $\nabla f_{i}(x)^{\top}$, this reads
$\operatorname{range}(JF(x))\cap(-\mathbb{R}_{++})^{m}=\varnothing,$
where $\mathbb{R}_{++}$ is the set of (strictly) positive real numbers.
This geometric condition is equivalent to the convex hull of the gradients containing the origin. The proximity to this state is quantified by the following gap function.
Definition 3.5 (Pareto stationarity gap).
The Pareto stationarity gap at $x\in\mathbb{R}^{d}$ is
$\mathcal{G}(x)\;:=\;\min_{\lambda\in\Delta^{m}}\Big\|\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)\Big\|,\qquad\Delta^{m}:=\big\{\lambda\in\mathbb{R}^{m}:\ \lambda_{i}\geq 0,\ \sum_{i=1}^{m}\lambda_{i}=1\big\}.$
We call $x$ $\varepsilon$-Pareto stationary if $\mathcal{G}(x)\leq\varepsilon$; the case $\varepsilon=0$ corresponds to Pareto stationarity.
Relation to Pareto merit functions
Besides stationarity gaps, the MOO literature also uses merit (value-gap) functions. Assume that each $f_{i}$ is convex and differentiable. For instance, consider the Pareto merit function
$u_{0}(x)\;:=\;\sup_{z\in\mathbb{R}^{d}}\min_{i=1,\dots,m}\bigl(f_{i}(x)-f_{i}(z)\bigr),$
which vanishes at weakly Pareto optima and is positive otherwise (see, e.g., [34]).
Let $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$ for $\lambda\in\Delta^{m}$, and assume $\arg\min f_{\lambda}\neq\emptyset$ for all $\lambda\in\Delta^{m}$; pick any minimizer $x_{\lambda}^{\star}\in\arg\min f_{\lambda}$ and write $f_{\lambda}^{\star}:=f_{\lambda}(x_{\lambda}^{\star})$. We can upper bound $u_{0}(x)$ by scalarization gaps in three simple steps. Fix $x$ and $z$ and apply the identity $\min_{i}a_{i}=\min_{\lambda\in\Delta^{m}}\sum_{i}\lambda_{i}a_{i}$ with $a_{i}=f_{i}(x)-f_{i}(z)$, to obtain
$\min_{i=1,\dots,m}\bigl(f_{i}(x)-f_{i}(z)\bigr)=\min_{\lambda\in\Delta^{m}}\sum_{i=1}^{m}\lambda_{i}\bigl(f_{i}(x)-f_{i}(z)\bigr)=\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Substituting into the definition of $u_{0}$ yields
$u_{0}(x)=\sup_{z\in\mathbb{R}^{d}}\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Next, we use the elementary min–max inequality
$\sup_{z}\min_{\lambda}\Phi(z,\lambda)\leq\min_{\lambda}\sup_{z}\Phi(z,\lambda)$
with $\Phi(z,\lambda)=f_{\lambda}(x)-f_{\lambda}(z)$ to obtain
$u_{0}(x)\;\leq\;\min_{\lambda\in\Delta^{m}}\sup_{z\in\mathbb{R}^{d}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr).$
Finally, for each fixed $\lambda$, since $f_{\lambda}(x)$ does not depend on $z$,
$\sup_{z\in\mathbb{R}^{d}}\bigl(f_{\lambda}(x)-f_{\lambda}(z)\bigr)=f_{\lambda}(x)-\inf_{z\in\mathbb{R}^{d}}f_{\lambda}(z)=f_{\lambda}(x)-f_{\lambda}^{\star},$
and therefore
$u_{0}(x)\;\leq\;\min_{\lambda\in\Delta^{m}}\bigl(f_{\lambda}(x)-f_{\lambda}^{\star}\bigr).$
By convexity, for each $\lambda$,
$f_{\lambda}(x)-f_{\lambda}^{\star}\leq\langle\nabla f_{\lambda}(x),\,x-x_{\lambda}^{\star}\rangle\leq\|\nabla f_{\lambda}(x)\|\,\|x-x_{\lambda}^{\star}\|.$
Therefore,
$u_{0}(x)\;\leq\;\bar{R}(x)\,\mathcal{G}(x),\qquad\bar{R}(x):=\sup_{\lambda\in\Delta^{m}}\|x-x_{\lambda}^{\star}\|.$
Moreover, if each $f_{i}$ is $\mu$-strongly convex, then so is every $f_{\lambda}$, and the Polyak–Łojasiewicz inequality yields $f_{\lambda}(x)-f_{\lambda}^{\star}\leq\frac{1}{2\mu}\|\nabla f_{\lambda}(x)\|^{2}$; hence
$u_{0}(x)\;\leq\;\frac{1}{2\mu}\,\mathcal{G}(x)^{2}.$
In particular, under a boundedness condition ensuring $\bar{R}(x)<\infty$ on the region of interest (or under strong convexity), any convergence guarantee stated in terms of the stationarity gap $\mathcal{G}(x)$ immediately implies a corresponding bound for the merit function $u_{0}(x)$.
Equivalence of stationarity notions
The next lemma makes precise that Pareto criticality is equivalent to Pareto stationarity $\mathcal{G}(x)=0$.
Lemma 3.6 (No common descent $\Longleftrightarrow$ convex-hull stationarity).
For $C^{1}$ objectives $(f_{i})$, the following are equivalent at a point $x$:
  1. (a)
    (Pareto criticality) There is no $v\in\mathbb{R}^{d}$ with $\nabla f_{i}(x)^{\top}v<0$ for all $i$.
  2. (b)
    (Convex-hull stationarity) There exists $\lambda\in\Delta^{m}$ such that
    $\sum_{i=1}^{m}\lambda_{i}\,\nabla f_{i}(x)=0$
    (equivalently, $\mathcal{G}(x)=0$).
Proof.
Apply Gordan’s theorem of the alternative to $A=JF(x)$: exactly one of $Av<0$ or $A^{\top}y=0$ with $y\geq 0$, $y\neq 0$ holds. The negation of the former yields a nonzero $y\geq 0$ with $JF(x)^{\top}y=0$; normalizing $y$ to the simplex gives (b). Conversely, (b) rules out any $v$ with $JF(x)v<0$ because $\sum_{i}\lambda_{i}\,\nabla f_{i}(x)^{\top}v$ would be a convex combination of strictly negative numbers. ∎
Necessary and sufficient conditions
We now record the basic relationships between optimality and stationarity in the unconstrained setting.
Proposition 3.7 (General $C^{1}$ case: necessity).
If $x^{\star}$ is weakly Pareto optimal (for not-necessarily convex $C^{1}$ objectives), then $x^{\star}$ is Pareto stationary; equivalently $\mathcal{G}(x^{\star})=0$.
Proof.
At a weakly Pareto optimal point there is no direction that strictly decreases all objectives. By Lemma 3.6, this is equivalent to the existence of $\lambda\in\Delta^{m}$ with $\sum_{i}\lambda_{i}\,\nabla f_{i}(x^{\star})=0$, hence $\mathcal{G}(x^{\star})=0$. ∎
Proposition 3.8 (Convex $C^{1}$ case: characterization of weak Pareto optima).
Suppose each $f_{i}$ is convex and $C^{1}$. Then, for $x^{\star}\in\mathbb{R}^{d}$, the following are equivalent:
  1. (i)
    $x^{\star}$ is weakly Pareto optimal;
  2. (ii)
    $x^{\star}$ is Pareto stationary, i.e., $\mathcal{G}(x^{\star})=0$;
  3. (iii)
    there exists $\lambda\in\Delta^{m}$ such that $x^{\star}\in\arg\min_{x}f_{\lambda}(x)$ with $f_{\lambda}:=\sum_{i=1}^{m}\lambda_{i}f_{i}$.
Proof.
[…]
Remark 3.9.
Under convexity, Proposition 3.8 shows that $\mathcal{G}(\cdot)$ is an exact first-order optimality measure for weak Pareto optimality: $\mathcal{G}(x^{\star})=0$ if and only if $x^{\star}$ is weakly Pareto optimal. In the general (nonconvex) $C^{1}$ case, $\mathcal{G}(x^{\star})=0$ remains a necessary condition but is not sufficient, as usual for first-order stationarity.
3.4 Oracle Model and Algorithm Classes
We formally define the classes of first-order algorithms considered in this work. The broadest class contains nearly all modern iterative methods.
Definition 3.10 (Span First-Order Methods).
An algorithm is a span first-order method if its iterates satisfy
$x^{(t)}\;\in\;x^{(0)}+\mathrm{span}\{\nabla f_{i}(x^{(j)}):\ i=1,\dots,m,\ j=0,\dots,t-1\}.$
This general class allows for adaptive step sizes and momentum terms that depend on the entire history of observed gradients.
Definition 3.11 (Oblivious Span First-Order Methods).
A span first-order method is oblivious if, for each $t\geq 1$, there exist coefficients $\{\beta^{(t)}_{i,j}\}_{i=1,\ldots,m;\,j=0,\ldots,t-1}$ that are fixed in advance (possibly as a function of known problem parameters such as $L,\mu,m,T$) and do not depend on the realized oracle replies, such that
$x^{(t)}\;=\;x^{(0)}+\sum_{j=0}^{t-1}\sum_{i=1}^{m}\beta^{(t)}_{i,j}\,\nabla f_{i}\bigl(x^{(j)}\bigr).$
Equivalently, the method satisfies the span property and uses only pre-scheduled (instance-independent) linear combination coefficients.
Remark 3.12.
The class of oblivious span methods includes accelerated schemes with fixed parameter schedules such as Nesterov’s accelerated gradient method on a fixed scalarization.
For our sharpest convex lower bound, we consider a more restricted, non-adaptive class.
Definition 3.13 (Oblivious One-Step Gradient Methods).
Fix $\lambda\in\Delta^{m}$ and define $f_{\lambda}(x)=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$. An oblivious one-step gradient method is an algorithm that generates iterates by
$x^{(t+1)}=x^{(t)}-\alpha_{t}\,\nabla f_{\lambda}(x^{(t)}),\qquad t=0,1,2,\dots,$
with pre-scheduled step sizes $\{\alpha_{t}\}$ satisfying $0\leq\alpha_{t}\leq 1/L$ for all $t$ that may depend only on known problem parameters (e.g., $L$, $\mu$), but not on the oracle’s feedback. For a quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$, the error satisfies $x^{(t)}-x^{\star}=p_{t}(H)(x^{(0)}-x^{\star})$ with
$p_{t}(\zeta)\;=\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\zeta\bigr),\qquad p_{t}(0)=1.$
Remark 3.14.
The oblivious one-step class models standard gradient descent with a fixed step schedule. It does not include methods with momentum, such as AGD or Polyak’s heavy-ball method.
Lemma 3.15 (Krylov Representation for Span Methods on Quadratics).
Let $g:\mathbb{R}^{d}\to\mathbb{R}$ be the quadratic
$g(x)=\tfrac{1}{2}\,x^{\top}Hx-b^{\top}x,$
where $H\in\mathbb{R}^{d\times d}$ is symmetric and $b\in\operatorname{range}(H)$, so that $g$ attains its minimum and there exists $x^{\star}\in\arg\min g$ with $Hx^{\star}=b$. Consider any iterative algorithm whose iterates $\{x^{(t)}\}_{t\geq 0}$ satisfy the span property
$x^{(t)}\in x^{(0)}+\operatorname{span}\{\nabla g(x^{(0)}),\ldots,\nabla g(x^{(t-1)})\}\quad\text{for all }t\geq 1.$ (2)
Then, for every $t\geq 0$, there exists a real polynomial $p_{t}$ of degree at most $t$ with $p_{t}(0)=1$ such that
$x^{(t)}-x^{\star}\;=\;p_{t}(H)\,\bigl(x^{(0)}-x^{\star}\bigr).$
Moreover, if the method is an oblivious span method (Definition 3.11) when applied to $g$, then for each $t$ the polynomial $p_{t}$ depends only on the method’s fixed coefficients (and on $t$), and is therefore independent of the specific quadratic instance $(H,b)$.
Furthermore, if the method is an oblivious one-step gradient method on $g$, i.e.,
$x^{(t+1)}=x^{(t)}-\alpha_{t}\,\nabla g(x^{(t)})\quad(t\geq 0)$
for some pre-scheduled stepsizes $\{\alpha_{t}\}_{t\geq 0}$, then
$p_{t}(\zeta)=\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr)\qquad(t\geq 1),\qquad p_{0}\equiv 1.$
The same conclusions hold when $g$ is any fixed scalarization of a quadratic multiobjective instance.
Proof.
Fix any minimizer $x^{\star}$ with $Hx^{\star}=b$; such a point exists by the assumption $b\in\operatorname{range}(H)$. For every $x$, we then have
$\nabla g(x)=Hx-b=H\,(x-x^{\star}).$ (3)
Let $e^{(t)}:=x^{(t)}-x^{\star}$ and $e^{(0)}:=x^{(0)}-x^{\star}$. We prove by induction on $t$ that
$e^{(t)}\;=\;p_{t}(H)\,e^{(0)}\quad\text{for some polynomial }p_{t}\text{ with }\deg p_{t}\leq t\text{ and }p_{t}(0)=1.$ (4)
Base case $t=0$. Trivially, $e^{(0)}=I\,e^{(0)}$, so (4) holds with $p_{0}\equiv 1$.
Inductive step. Assume (4) holds for all indices up to $t$. By the span property (2), there exist scalars $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ such that
$x^{(t+1)}\;=\;x^{(0)}+\sum_{j=0}^{t}\beta_{j}^{(t)}\,\nabla g\bigl(x^{(j)}\bigr).$
(If the method is oblivious span, we take $\{\beta_{j}^{(t)}\}$ to be exactly the pre-scheduled coefficients specified by the method.) Subtracting $x^{\star}$ and using (3) gives
$e^{(t+1)}\;=\;e^{(0)}+\sum_{j=0}^{t}\beta_{j}^{(t)}\,H\,e^{(j)}.$ (5)
By the induction hypothesis, for each $0\leq j\leq t$ there exists a polynomial $p_{j}$ with $\deg p_{j}\leq j$ and $p_{j}(0)=1$ such that $e^{(j)}=p_{j}(H)\,e^{(0)}$. Substituting into (5) yields
$e^{(t+1)}\;=\;\Bigl(I+H\,\sum_{j=0}^{t}\beta_{j}^{(t)}\,p_{j}(H)\Bigr)\,e^{(0)}.$
Define the polynomial
$q_{t}(\zeta)\;:=\;\sum_{j=0}^{t}\beta_{j}^{(t)}\,p_{j}(\zeta),\qquad\text{so that}\qquad\deg q_{t}\leq\max_{0\leq j\leq t}\deg p_{j}\leq t,$
and set
$p_{t+1}(\zeta)\;:=\;1+\zeta\,q_{t}(\zeta).$
Then $\deg p_{t+1}\leq t+1$, $p_{t+1}(0)=1$, and
$e^{(t+1)}\;=\;p_{t+1}(H)\,e^{(0)}.$
This completes the induction and proves (4) for all $t\geq 0$.
Obliviousness and instance-independence. If the method is an oblivious span method (Definition 3.11), then the coefficients $\{\beta_{j}^{(t)}\}_{j=0}^{t}$ in (5) are fixed in advance (they may depend on $t$ and on known problem parameters, but not on the realized oracle replies). Hence the recursion that constructs $p_{t+1}$ from $(p_{0},\ldots,p_{t})$ uses only these fixed coefficients. Therefore, for each $t$, the resulting polynomial $p_{t}$ is determined solely by the method’s fixed coefficients (and by $t$), and in particular does not depend on $(H,b)$.
It remains to establish the stated product form for oblivious one-step methods. In that case,
$e^{(t+1)}\;=\;x^{(t)}-x^{\star}-\alpha_{t}\,\nabla g(x^{(t)})\;=\;\bigl(I-\alpha_{t}H\bigr)\,e^{(t)}\qquad(t\geq 0),$
using (3). Iterating this linear recurrence and recalling $e^{(0)}=p_{0}(H)\,e^{(0)}$ with $p_{0}\equiv 1$ gives
$e^{(t)}\;=\;\Bigl(\prod_{k=0}^{t-1}\bigl(I-\alpha_{k}H\bigr)\Bigr)\,e^{(0)}\;=\;p_{t}(H)\,e^{(0)},$
with $p_{t}(\zeta)=\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)$, which has degree $t$ and satisfies $p_{t}(0)=1$.
Independence of the choice of minimizer. If $H$ is singular, the minimizer may not be unique: any $x^{\star}$ with $Hx^{\star}=b$ differs from another minimizer by a vector in $\ker(H)$. Let $x^{\star}{}^{\prime}$ be another minimizer and write $x^{\star}{}^{\prime}=x^{\star}+z$ with $Hz=0$. Then $e^{\prime(0)}=x^{(0)}-x^{\star}{}^{\prime}=e^{(0)}-z$ and, by the proven representation,
$x^{(t)}-x^{\star}{}^{\prime}\;=\;p_{t}(H)\,e^{\prime(0)}\;=\;p_{t}(H)\,e^{(0)}-p_{t}(0)\,z.$
Since $x^{(t)}-x^{\star}{}^{\prime}=(x^{(t)}-x^{\star})-z$, consistency for all $z\in\ker(H)$ requires and is ensured by $p_{t}(0)=1$. Thus the representation is well defined independently of the chosen minimizer. ∎
4 Oracle Complexity Lower Bounds
4.1 A Non-Degenerate Lifting Construction
A common concern with oracle lower bounds is that the hard instances may be degenerate from a multiobjective perspective—for example, the objectives may coincide up to constants, or the Pareto set may collapse to a singleton—so that the lower bound reflects a scalar artifact rather than a genuinely multiobjective phenomenon. To rule this out, we introduce a simple but robust lifting device: it embeds an arbitrary scalar function $g$ defined on a subspace $V$ into an $m$-objective instance on $\mathbb{R}^{d}=V\oplus W$ by adding objective-dependent quadratic terms on a complementary subspace $W$. The resulting objectives are distinct, the Pareto set is non-singleton, and the Pareto stationarity gap $\mathcal{G}$ admits an explicit decomposition showing that it dominates the scalar stationarity measure $\|\nabla g\|$. This “gap domination” property is the mechanism that transfers first-order lower bounds from the scalar problem to the multiobjective setting (including nonconvex $g$, since convexity is not used in the domination argument).
Theorem 4.1 (Non-degenerate lifting and stationarity-gap domination).
Fix integers $m\geq 2$ and $d\geq 1$. Let $\mathbb{R}^{d}=V\oplus W$ be an orthogonal direct sum with $\dim(W)\geq m-1$, and pick $m$ affinely independent points $a_{1},\dots,a_{m}\in W$. Let $g:V\to\mathbb{R}$ be $C^{1}$ with $L_{g}$-Lipschitz gradient, and fix $\gamma>0$. Define for $i=1,\dots,m$:
$f_{i}(x):=g(x_{V})+\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2},\qquad x=x_{V}+x_{W}.$ (6)
Then:
  1. (i)
    The objectives $\{f_{i}\}_{i=1}^{m}$ are distinct.
  2. (ii)
    Each $f_{i}$ is $L$-smooth with $L=\max\{L_{g},\gamma\}$. If $g$ is convex (resp. $\mu$-strongly convex) and $\gamma\geq 0$ (resp. $\gamma\geq\mu$), then each $f_{i}$ is convex (resp. $\mu$-strongly convex).
  3. (iii)
    For all $x\in\mathbb{R}^{d}$,
    $\mathcal{G}(x)^{2}=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\,\mathrm{dist}\!\left(x_{W},\mathrm{conv}\{a_{1},\dots,a_{m}\}\right)^{2},$ (7)
    in particular $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
  4. (iv)
    If $g$ is convex and admits a minimizer (i.e., $\arg\min g\neq\emptyset$), then the (weak) Pareto set equals
    $\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
    In particular, $\mathrm{conv}\{a_{1},\dots,a_{m}\}$ is a non-singleton polytope. If $g$ is $\mu$-strongly convex, then $\arg\min g=\{x_{V}^{\star}\}$ is a singleton and $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}$.
  5. (v)
    Regardless of convexity of $g$, the Pareto-stationary set of the lifted instance satisfies
    $\mathcal{S}:=\{x:\mathcal{G}(x)=0\}=\{x_{V}\in V:\nabla g(x_{V})=0\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Consequently, any oracle lower bound on $\|\nabla g(u^{(t)})\|$ for a scalar first-order method applied to $g$ transfers verbatim to a lower bound on $\mathcal{G}(x^{(t)})$ for the lifted MOO instance.
Proof.
Throughout, we use the orthogonal decomposition $\mathbb{R}^{d}=V\oplus W$ and write $x=x_{V}+x_{W}$ with $x_{V}\in V$ and $x_{W}\in W$. Differentiating (6) yields, for each $i$,
$\nabla f_{i}(x)=\bigl(\nabla g(x_{V}),\ \gamma(x_{W}-a_{i})\bigr).$ (8)
(i) Distinctness of objectives. Because the points $a_{1},\ldots,a_{m}$ are affinely independent, in particular they are pairwise distinct. Fix $i\neq j$ and take any $x_{V}\in V$ with $x_{W}=a_{i}$. Then
$f_{i}(x_{V}+a_{i})=g(x_{V}),\qquad f_{j}(x_{V}+a_{i})=g(x_{V})+\frac{\gamma}{2}\|a_{i}-a_{j}\|^{2}>g(x_{V}),$
so $f_{i}\not\equiv f_{j}$. Hence the objectives are distinct.
(ii) Smoothness and (strong) convexity. We first verify smoothness. Since $g$ has $L_{g}$-Lipschitz gradient on $V$, for all $u,v\in V$, $\|\nabla g(u)-\nabla g(v)\|\leq L_{g}\|u-v\|$. Using (8), we obtain for any $x,y\in\mathbb{R}^{d}$,
$\|\nabla f_{i}(x)-\nabla f_{i}(y)\|^{2}$ $=\|\nabla g(x_{V})-\nabla g(y_{V})\|^{2}+\gamma^{2}\|x_{W}-y_{W}\|^{2}$
$\leq L_{g}^{2}\|x_{V}-y_{V}\|^{2}+\gamma^{2}\|x_{W}-y_{W}\|^{2}$
$\leq\max\{L_{g}^{2},\gamma^{2}\}\bigl(\|x_{V}-y_{V}\|^{2}+\|x_{W}-y_{W}\|^{2}\bigr)$
$=\max\{L_{g}^{2},\gamma^{2}\}\,\|x-y\|^{2},$
where we used orthogonality of $V$ and $W$ in the last step. Taking square roots yields $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq\max\{L_{g},\gamma\}\|x-y\|$, hence each $f_{i}$ is $L$-smooth with $L=\max\{L_{g},\gamma\}$.
If $g$ is convex, then $x\mapsto g(x_{V})$ is convex on $\mathbb{R}^{d}$ (as a composition of the linear projection $x\mapsto x_{V}$ with a convex function), and $x\mapsto\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2}$ is convex for $\gamma\geq 0$; therefore $f_{i}$ is convex. If $g$ is $\mu$-strongly convex on $V$, then $x\mapsto g(x_{V})$ is $\mu$-strongly convex along directions in $V$, while $x\mapsto\frac{\gamma}{2}\|x_{W}-a_{i}\|^{2}$ is $\gamma$-strongly convex along directions in $W$; consequently, for all $x,y\in\mathbb{R}^{d}$,
$f_{i}(y)\geq f_{i}(x)+\langle\nabla f_{i}(x),y-x\rangle+\frac{\min\{\mu,\gamma\}}{2}\|y-x\|^{2},$
so $f_{i}$ is $\min\{\mu,\gamma\}$-strongly convex. In particular, if $\gamma\geq\mu$ then each $f_{i}$ is $\mu$-strongly convex.
(iii) Stationarity-gap formula and domination. For any $\lambda\in\Delta^{m}$, summing (8) gives
$\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)=\left(\nabla g(x_{V}),\ \gamma\Bigl(x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\Bigr)\right).$
By orthogonality of $V$ and $W$,
$\left\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)\right\|^{2}=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\left\|x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\right\|^{2}.$
The first term is independent of $\lambda$, and $\{\sum_{i}\lambda_{i}a_{i}:\lambda\in\Delta^{m}\}=\mathrm{conv}\{a_{1},\ldots,a_{m}\}$, hence
$\mathcal{G}(x)^{2}$ $=\min_{\lambda\in\Delta^{m}}\left\|\sum_{i=1}^{m}\lambda_{i}\nabla f_{i}(x)\right\|^{2}$
$=\|\nabla g(x_{V})\|^{2}+\gamma^{2}\,\mathrm{dist}\!\left(x_{W},\mathrm{conv}\{a_{1},\ldots,a_{m}\}\right)^{2},$
which is (7). In particular, $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
(iv) Pareto set in the convex case. Assume $g$ is convex and admits a minimizer $x_{V}^{\star}\in\arg\min g$. By Proposition 3.8, weak Pareto optima coincide with minimizers of some scalarization $f_{\lambda}=\sum_{i=1}^{m}\lambda_{i}f_{i}$. For any $\lambda\in\Delta^{m}$,
$f_{\lambda}(x_{V},x_{W})$ $=g(x_{V})+\frac{\gamma}{2}\sum_{i=1}^{m}\lambda_{i}\|x_{W}-a_{i}\|^{2}$
$=g(x_{V})+\frac{\gamma}{2}\left\|x_{W}-\bar{a}(\lambda)\right\|^{2}+\mathrm{const}(\lambda),\qquad\bar{a}(\lambda):=\sum_{i=1}^{m}\lambda_{i}a_{i},$
where $\mathrm{const}(\lambda)$ is independent of $(x_{V},x_{W})$. Thus minimization decouples: the $V$-component is minimized by any $x_{V}\in\arg\min g$, and the quadratic in $x_{W}$ is uniquely minimized at $x_{W}=\bar{a}(\lambda)$. Hence
$\arg\min_{x\in\mathbb{R}^{d}}f_{\lambda}(x)=(\arg\min g)\times\{\bar{a}(\lambda)\}.$
As $\lambda$ ranges over $\Delta^{m}$, $\bar{a}(\lambda)$ ranges over $\mathrm{conv}\{a_{1},\ldots,a_{m}\}$, yielding
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Since $m\geq 2$ and the $a_{i}$ are affinely independent, $\mathrm{conv}\{a_{i}\}$ is a non-singleton polytope. If $g$ is $\mu$-strongly convex, then $\arg\min g=\{x_{V}^{\star}\}$ is a singleton and the same description reduces to $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}$.
(v) Pareto-stationary set for general $g$. By (7), $\mathcal{G}(x)=0$ holds if and only if $\nabla g(x_{V})=0$ and $\mathrm{dist}(x_{W},\mathrm{conv}\{a_{i}\})=0$, i.e., $x_{W}\in\mathrm{conv}\{a_{1},\ldots,a_{m}\}$. Therefore,
$\mathcal{S}=\{x:\mathcal{G}(x)=0\}=\{x_{V}\in V:\nabla g(x_{V})=0\}\times\mathrm{conv}\{a_{1},\ldots,a_{m}\}.$
Transfer claim. The pointwise domination $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ from (iii) implies that for any iterate sequence $(x^{(t)})_{t\geq 0}$,
$\mathcal{G}(x^{(t)})\geq\|\nabla g(x_{V}^{(t)})\|\qquad\text{for all }t.$
Hence any lower bound on $\|\nabla g(\cdot)\|$ along the projected iterates transfers immediately to a lower bound on $\mathcal{G}(\cdot)$ along the lifted iterates. ∎
4.2 Scalar hard instances: polynomial-degree and geometric hardness
We will use two types of scalar lower bounds: (i) quadratic lower bounds for span/oblivious methods that reduce to extremal polynomial problems; and (ii) geometric (zero-chain) lower bounds for finding stationary points under a bounded initial value gap, due to [6, 7].
4.2.1 Quadratic hard instances for span/oblivious methods
We first construct scalar quadratic hard instances tailored to span/oblivious first-order methods. For quadratics, these methods generate iterates that admit a polynomial/Krylov representation, which enables explicit finite-horizon lower bounds on $\min_{t\leq T}\|\nabla g(x^{(t)})\|$ and separates the behavior of the different method classes. Lemma 4.2 below states the specific scalar quadratic instances and bounds that will serve as base ingredients for our multiobjective lower bounds via the lifting theorem.
Lemma 4.2 (Hard scalar quadratic instances).
Fix $T\in\mathbb{N}$ and $R>0$.
  1. (a)
    (Strongly convex case) Fix $L\geq\mu>0$ and $\kappa:=L/\mu$. For any span first-order method making $T$ oracle calls on quadratics, there exist a dimension $d\geq T+1$, an $L$-smooth $\mu$-strongly convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$ such that the $T$-th iterate satisfies
    $\|\nabla g(x^{(T)})\|\;\geq\;\mu\,R\,\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
  2. (b)
    (Convex, oblivious one-step) Fix $L>0$. For any oblivious one-step gradient method run for $T$ steps with step sizes $\alpha_{t}\in[0,1/L]$, there exists an $L$-smooth convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$, where $x^{\star}\in\arg\min g$ is chosen as a closest minimizer to $x^{(0)}$, such that
    $\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\frac{LR}{4(T+1)}.$
  3. (c)
    (Convex, oblivious span: polynomial-degree) Fix $L>0$. For any oblivious span first-order method run for $T$ steps, there exists an $L$-smooth convex quadratic $g$ and an initialization with $\|x^{(0)}-x^{\star}\|=R$, where $x^{\star}\in\arg\min g$ is chosen as a closest minimizer to $x^{(0)}$, such that
    $\|\nabla g(x^{(T)})\|\;\geq\;\frac{LR}{2(T+1)^{2}}.$
Proof.
We prove each part separately.
(a) Strongly convex case. Let $g(x)=\frac{1}{2}x^{\top}Hx-b^{\top}x$ with $H\succ 0$. For any iterate sequence $\{x^{(t)}\}$ produced by a span first-order method on a quadratic, the error admits the polynomial representation (Lemma 3.15):
$x^{(t)}-x^{\star}\;=\;p_{t}(H)\,\bigl(x^{(0)}-x^{\star}\bigr),\qquad\deg p_{t}\leq t,\quad p_{t}(0)=1.$ (9)
Set $e^{(0)}:=x^{(0)}-x^{\star}$ with $\|e^{(0)}\|=R$. Since $H\succeq\mu I$, we have for any vector $v$ that $\|Hv\|\geq\mu\|v\|$, hence
$\bigl\|\nabla g(x^{(T)})\bigr\|=\bigl\|H\,p_{T}(H)\,e^{(0)}\bigr\|\;\geq\;\mu\,\bigl\|p_{T}(H)\,e^{(0)}\bigr\|.$ (10)
We now take worst case over the instance $(H,e^{(0)})$ with $\sigma(H)\subset[\mu,L]$ and $\|e^{(0)}\|=R$. For any fixed degree-$T$ polynomial $p_{T}$ with $p_{T}(0)=1$,
$\sup_{\begin{subarray}{c}\sigma(H)\subset[\mu,L]\\ \|e^{(0)}\|=R\end{subarray}}\bigl\|p_{T}(H)\,e^{(0)}\bigr\|\;=\;R\;\sup_{\zeta\in[\mu,L]}|p_{T}(\zeta)|,$
because for $H$ symmetric the operator norm $\|p_{T}(H)\|$ equals $\max_{\zeta\in\sigma(H)}|p_{T}(\zeta)|$, and we may align $e^{(0)}$ with an eigenvector where the maximum is attained. Therefore, combining with (10) and then minimizing over the (algorithm-induced) polynomial $p_{T}$ yields
$\sup_{\begin{subarray}{c}\sigma(H)\subset[\mu,L],\,\|e^{(0)}\|=R\end{subarray}}\bigl\|\nabla g(x^{(T)})\bigr\|\;\geq\;\mu R\cdot\inf_{\begin{subarray}{c}\deg p_{T}\leq T\\ p_{T}(0)=1\end{subarray}}\;\max_{\zeta\in[\mu,L]}|p_{T}(\zeta)|.$ (11)
The minimax problem on the right-hand side is the classical Chebyshev extremal problem on an interval with an off-interval normalization (at $\zeta=0$). The solution is detailed in Appendix A.3. Let
$\xi\;=\;\frac{2\zeta-(L+\mu)}{L-\mu}\in[-1,1],\qquad\xi_{0}\;=\;\frac{-\,(L+\mu)}{L-\mu}\;=\;-\,\frac{\kappa+1}{\kappa-1},$
and define $\tilde{p}_{T}(\xi):=p_{T}(\zeta(\xi))$. Then the extremal value is achieved by the scaled Chebyshev polynomial of the first kind,
$\tilde{p}_{T}^{\star}(\xi)\;=\;\frac{T_{T}(\xi)}{T_{T}(\xi_{0})},\qquad\min_{\deg\tilde{p}_{T}\leq T,\;\tilde{p}_{T}(\xi_{0})=1}\;\max_{\xi\in[-1,1]}|\tilde{p}_{T}(\xi)|\;=\;\frac{1}{|T_{T}(\xi_{0})|}.$
For $|x|>1$ one has $T_{T}(x)=\tfrac{1}{2}\bigl(\rho^{T}+\rho^{-T}\bigr)$ with $\rho=x+\sqrt{x^{2}-1}$. Substituting $x=|\xi_{0}|=\frac{\kappa+1}{\kappa-1}$ gives
$|T_{T}(\xi_{0})|\;=\;\frac{\rho^{T}+\rho^{-T}}{2},\qquad\rho=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
Thus the extremal value of (11) equals $\mu R\cdot\frac{2}{\rho^{T}+\rho^{-T}}$, proving
$\sup_{\text{instances}}\bigl\|\nabla g(x^{(T)})\bigr\|\;\geq\;\mu R\cdot\frac{2}{\rho^{T}+\rho^{-T}}\;\geq\;\mu R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},$
where the last inequality follows from $\frac{2}{a+a^{-1}}\geq a^{-1}$ for $a\geq 1$.
(b) Convex oblivious one-step case. This is precisely Corollary A.1.4 in Appendix A.1 (with $t=T$).
(c) Convex oblivious span polynomial-degree case. Fix an arbitrary oblivious span method and consider any $L$-smooth convex quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ with $\sigma(H)\subset[0,L]$. By Lemma 3.15 (oblivious case), there exists a polynomial $p_{T}$ with $\deg p_{T}\leq T$ and $p_{T}(0)=1$ that depends only on the method (and on $T$), such that
$x^{(T)}-x^{\star}=p_{T}(H)\,(x^{(0)}-x^{\star}),$
and hence
$\nabla g(x^{(T)})=H\,p_{T}(H)\,(x^{(0)}-x^{\star}).$
Let $\|x^{(0)}-x^{\star}\|=R$. For any fixed $\zeta\in[0,L]$, choose $H$ diagonal with a single eigenvalue $\zeta$ (and all remaining eigenvalues equal to 0) and align $x^{(0)}-x^{\star}$ with the corresponding eigenvector. Then
$\|\nabla g(x^{(T)})\|\;=\;R\,|\zeta\,p_{T}(\zeta)|.$
Since $p_{T}$ is fixed independently of $\zeta$ (obliviousness), taking the worst case over $\zeta\in[0,L]$ yields
$\sup_{\begin{subarray}{c}g\ \text{$L$-smooth convex quadratic}\\ \|x^{(0)}-x^{\star}\|=R\end{subarray}}\|\nabla g(x^{(T)})\|\;\geq\;R\,\max_{\zeta\in[0,L]}|\zeta p_{T}(\zeta)|.$
Finally, taking the infimum over oblivious span methods is lower bounded by the infimum over all admissible polynomials, and Appendix A.2 gives
$\inf_{\begin{subarray}{c}\deg p_{T}\leq T\\ p_{T}(0)=1\end{subarray}}\ \max_{\zeta\in[0,L]}|\zeta p_{T}(\zeta)|\;\geq\;\frac{L}{2(T+1)^{2}}.$
This yields $\|\nabla g(x^{(T)})\|\geq\frac{LR}{2(T+1)^{2}}$ as claimed. ∎
4.2.2 Geometric stationary-point lower bounds under value-gap bounds
We next state the geometric scalar stationary-point lower bounds that drive our new convex and nonconvex MOO results.
Lemma 4.3 (Convex $L$-smooth stationary-point lower bound under bounded value gap [7, Theorem 1]).
Fix the initial point $x^{(0)}=0$. A deterministic first-order method $\mathcal{A}$ is any procedure that, for $t=0,1,2,\dots$, adaptively chooses query points
$x^{(t+1)}=\Phi_{t}\!\left(\{(g(x^{(s)}),\nabla g(x^{(s)}))\}_{s=0}^{t}\right),$
for some deterministic mappings $\Phi_{t}$, where the oracle returns $(g(x),\nabla g(x))$ at each query.
Fix $L>0$ and $\Delta>0$. For any deterministic first-order method $\mathcal{A}$ and any $T\in\mathbb{N}$, there exist a dimension $d\geq 2T$ and a convex function $g:\mathbb{R}^{d}\to\mathbb{R}$ with $L$-Lipschitz gradient such that
$g(x^{(0)})-\inf_{x\in\mathbb{R}^{d}}g(x)\leq\Delta,$
and the iterates $\{x^{(t)}\}_{t=0}^{T}$ generated by $\mathcal{A}$ on $g$ satisfy
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
Equivalently, any deterministic first-order method requires $T=\Omega\!\left(\sqrt{L\Delta}/\varepsilon\right)$ oracle calls in the worst case to guarantee an output $x$ with $\|\nabla g(x)\|\leq\varepsilon$ over this function class.
Proof.
Carmon, Duchi, Hinder, and Sidford [7, Theorem 1] prove the following: for any $\varepsilon>0$ and any deterministic first-order method $\mathcal{A}$, there exists a convex function $f$ with $L$-Lipschitz gradient and $f(x^{(0)})-\inf_{x}f(x)\leq\Delta$ such that $\mathcal{A}$ must make at least $\frac{\sqrt{L\Delta}}{4\varepsilon}$ first-order oracle calls to guarantee finding a point $x$ with $\|\nabla f(x)\|\leq\varepsilon$. (Their statement is given for an orthogonally invariant function class $K_{1}(\Delta,L)$, and the reduction from zero-respecting to general deterministic methods is handled via a resisting-oracle / rotation argument, which increases the dimension to at most $2T$.)
Now fix $T\in\mathbb{N}$ and set $\varepsilon:=\frac{\sqrt{L\Delta}}{4(T+1)}$. Then $\frac{\sqrt{L\Delta}}{4\varepsilon}=T+1$, so any method using only $T$ oracle calls cannot guarantee $\|\nabla g(x)\|\leq\varepsilon$ on the above convex $L$-smooth class. Therefore, for the corresponding worst-case instance $g$, all iterates up to time $T$ satisfy $\|\nabla g(x^{(t)})\|>\varepsilon$, and hence
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\geq\varepsilon=\frac{\sqrt{L\Delta}}{4(T+1)}.$
Rearranging yields the equivalent complexity form $T=\Omega(\sqrt{L\Delta}/\varepsilon)$. ∎
Lemma 4.4 (Nonconvex $L$-smooth stationary-point lower bound (Lipschitz gradient only) [6, Theorem 1]).
Fix the initial point $x^{(0)}=0$. A deterministic first-order method $\mathcal{A}$ is any procedure that, for $t=0,1,2,\dots$, adaptively chooses query points
$x^{(t+1)}=\Phi_{t}\!\left(\{(g(x^{(s)}),\nabla g(x^{(s)}))\}_{s=0}^{t}\right),$
for some deterministic mappings $\Phi_{t}$, where the oracle returns the pair $(g(x),\nabla g(x))$ at each query.
There exists a universal constant $c>0$ such that for any $L>0$, $\Delta>0$, any deterministic first-order method $\mathcal{A}$, and any $T\in\mathbb{N}$, there exist a dimension $d\geq 2T+1$ and an $L$-smooth (not necessarily convex) function $g:\mathbb{R}^{d}\to\mathbb{R}$ satisfying
$g(x^{(0)})-\inf_{x\in\mathbb{R}^{d}}g(x)\leq\Delta,$
such that the iterates $\{x^{(t)}\}_{t=0}^{T}$ generated by $\mathcal{A}$ on $g$ obey
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;c\,\sqrt{\frac{L\Delta}{T+1}}.$
Equivalently, any deterministic first-order method requires $T=\Omega(L\Delta/\varepsilon^{2})$ oracle calls in the worst case to guarantee an output $x$ with $\|\nabla g(x)\|\leq\varepsilon$.
Proof.
Carmon, Duchi, Hinder, and Sidford prove (see [6, Theorem 1] specialized to $p=1$) that there exists a universal constant $C>0$ such that for every $\varepsilon>0$, every deterministic first-order method $\mathcal{A}$ requires at least
$T_{\varepsilon}\;\geq\;C\,\frac{L\Delta}{\varepsilon^{2}}$
oracle calls in the worst case over the class of $L$-smooth functions with initial gap at most $\Delta$, i.e., there exists an $L$-smooth (possibly nonconvex) $g$ with $g(x^{(0)})-\inf g\leq\Delta$ such that $\|\nabla g(x^{(t)})\|>\varepsilon$ for all $t<T_{\varepsilon}$.
Now fix $T\in\mathbb{N}$ and set $\varepsilon:=\sqrt{\frac{CL\Delta}{T+1}}$. Then $T+1<C\frac{L\Delta}{\varepsilon^{2}}$, hence $T<T_{\varepsilon}$ for the corresponding worst-case instance $g$, which implies $\|\nabla g(x^{(t)})\|>\varepsilon$ for all $t\leq T$ and therefore
$\min_{0\leq t\leq T}\|\nabla g(x^{(t)})\|\;\geq\;\varepsilon\;=\;\sqrt{\frac{CL\Delta}{T+1}}.$
Renaming $c:=\sqrt{C}$ yields the stated bound. The equivalence to $T=\Omega(L\Delta/\varepsilon^{2})$ follows by rearranging the inequality. ∎
4.3 Strongly Convex MOO: A Tight Linear Rate
We now state the corresponding multiobjective lower bound in the strongly convex regime.
Theorem 4.5 (Strongly Convex Lower Bound).
Fix $L\geq\mu>0$, $T\in\mathbb{N}$, $\kappa:=L/\mu$, and $R>0$. Consider any span first-order method that makes $T$ first-order oracle calls when applied to a quadratic instance. Then there exist integers $m\geq 2$ and $d\geq(T+1)+(m-1)$, an MOO instance $(f_{1},\dots,f_{m}):\mathbb{R}^{d}\to\mathbb{R}^{m}$ with $m$ objectives each $L$-smooth and $\mu$-strongly convex, and an initialization $x^{(0)}$ with $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the $T$-th iterate produced by the method satisfies
$\mathcal{G}\bigl(x^{(T)}\bigr)$ $\;\geq\;\mu\,R\;\frac{2}{\rho^{T}+\rho^{-T}}$
$\;\geq\;\mu\,R\left(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\right)^{T},\qquad\rho:=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
In particular, any such method has worst-case iteration complexity
$\Omega\!\bigl(\sqrt{\kappa}\log(1/\varepsilon)\bigr)$
to reduce the Pareto stationarity gap below $\varepsilon$.
Proof.
By Lemma 4.2(a), there exists a dimension $\dim(V)\geq T+1$ and an $L$-smooth $\mu$-strongly convex quadratic $g$ on a subspace $V$ that forces the stated lower bound on $\|\nabla g(u^{(T)})\|$ for any span method (with $\|u^{(0)}-u^{\star}\|=R$). Apply the lifting construction of Theorem 4.1 with $\gamma=\mu$ to obtain an MOO instance $(f_{1},\dots,f_{m})$ with $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ and
$\mathcal{G}(x^{(T)})\;\geq\;\|\nabla g(x_{V}^{(T)})\|.$
Since the $V$-component of the iterate sequence produced by a span method on the lifted instance satisfies the span property with respect to $g$, the scalar lower bound transfers, proving the claim. The iteration-complexity statement follows by solving $\mu R\bigl((\sqrt{\kappa}-1)/(\sqrt{\kappa}+1)\bigr)^{T}\leq\varepsilon$ for $T$. ∎
4.4 Convex MOO: Oblivious vs. Adaptive Methods
4.4.1 Oblivious One-Step Gradient Methods: An $\Omega(1/T)$ Min-Iterate Bound
For the more restricted class of oblivious one-step gradient methods, we establish a lower bound of order $\Omega(1/T)$. This bound applies to the minimum Pareto stationarity gap achieved over the first $T$ iterates and demonstrates a fundamental performance limit for non-adaptive algorithms that do not incorporate momentum or history-dependent updates.
Theorem 4.6 (Oblivious Convex Lower Bound).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any oblivious one-step gradient method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\frac{LR}{4(T+1)}.$
Proof.
By Lemma 4.2(b), for the given stepsize schedule there exists an $L$-smooth convex quadratic $g$ and an initialization $u^{(0)}$ such that
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{LR}{4(T+1)}\qquad\text{and}\qquad\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Choose $u^{\star}\in\arg\min g$ to be a closest minimizer to $u^{(0)}$, so that $\|u^{(0)}-u^{\star}\|=\mathrm{dist}(u^{(0)},\arg\min g)=R$.
Apply Theorem 4.1 with $\gamma=L$ to lift $g$ into a non-degenerate MOO instance $(f_{1},\dots,f_{m})$. By Theorem 4.1(iv), the (weak) Pareto set satisfies
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Pick $w^{(0)}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$ (e.g., $w^{(0)}=a_{1}$) and set $x^{(0)}=(u^{(0)},w^{(0)})\in V\oplus W$. Then
$\mathrm{dist}\!\bigl(x^{(0)},\mathcal{P}\bigr)=\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Finally, Theorem 4.1(iii) gives $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for all $x$. Hence,
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\min_{0\leq t\leq T}\|\nabla g(x_{V}^{(t)})\|\;\geq\;\frac{LR}{4(T+1)},$
as claimed. ∎
4.4.2 Oblivious span methods: polynomial-degree vs. geometric lower bounds
For oblivious span methods, the error polynomial $p_{t}$ is no longer restricted to have the product form of the one-step class, but it remains instance-independent. This leads to a universal $1/T^{2}$ last-iterate lower bound based on Markov’s inequality for polynomials. This bound is independent of geometric “chain” constructions and isolates a limitation coming solely from polynomial degree constraints within the oblivious span model.
For fully adaptive first-order methods (allowed to use the entire oracle transcript), we will also show a stronger geometric $\Omega(1/T)$ min-iterate lower bound under a bounded scalarization gap by lifting the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [7].
Theorem 4.7 (Polynomial-degree lower bound for oblivious span methods (last iterate)).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any oblivious span first-order method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the last iterate obeys
$\mathcal{G}(x^{(T)})\;\geq\;\frac{LR}{2\,(T+1)^{2}}.$
Proof.
By Lemma 4.2(c), there exists an $L$-smooth convex quadratic $g$ and an initialization $u^{(0)}$ such that for any oblivious span method run for $T$ steps,
$\|\nabla g(u^{(T)})\|\;\geq\;\frac{LR}{2\,(T+1)^{2}}\qquad\text{and}\qquad\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Choose $u^{\star}\in\arg\min g$ to be a closest minimizer to $u^{(0)}$, so that $\|u^{(0)}-u^{\star}\|=\mathrm{dist}(u^{(0)},\arg\min g)=R$.
Apply Theorem 4.1 with $\gamma=L$ to lift $g$ into a non-degenerate MOO instance $(f_{1},\dots,f_{m})$. By Theorem 4.1(iv),
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Pick $w^{(0)}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$ (e.g., $w^{(0)}=a_{1}$) and set $x^{(0)}=(u^{(0)},w^{(0)})\in V\oplus W$. Then
$\mathrm{dist}\!\bigl(x^{(0)},\mathcal{P}\bigr)=\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Finally, Theorem 4.1(iii) gives $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for all $x$. Therefore,
$\mathcal{G}(x^{(T)})\;\geq\;\|\nabla g(x_{V}^{(T)})\|\;\geq\;\frac{LR}{2\,(T+1)^{2}},$
as claimed. ∎
Remark 4.8.
Iterate selection. Theorem 4.6 is a min-iterate lower bound, whereas Theorem 4.7 is a last-iterate bound; the two statements are therefore not directly comparable a priori.
Remark 4.9 (The polynomial-degree bound is loose).
Theorem 4.7 is a last-iterate bound derived from polynomial extremal inequalities and does not exploit geometric hard instances (e.g., zero-chains). As we show next, geometric stationary-point lower bounds imply a stronger min-iterate $\Omega(1/T)$ lower bound for adaptive methods in smooth convex MOO under a bounded scalarization gap.
We now lift the convex stationary-point lower bound of Carmon–Duchi–Hinder–Sidford [7]. The appropriate “value gap” assumption in MOO is expressed through a scalarization $f_{\lambda}$.
Our lifting construction (Theorem 4.1) produces a multiobjective instance $f_{1},\dots,f_{m}$ from a scalar function $g$ such that the Pareto stationarity gap dominates the scalar gradient norm, i.e., $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for every $x=(x_{V},x_{W})$. For lower bounds proved for restricted algorithm classes (e.g., oblivious one-step or (oblivious) span methods), this domination property is sufficient: the iterates must remain in prescribed linear spans / polynomial images of past first-order information, so the $V$-component of the iterates on the lifted instance inherits the same restriction, and the corresponding scalar lower bounds transfer directly via $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
In contrast, the lower bounds under a bounded initial value gap $\Delta$ (such as those of [7]) are minimax statements over all deterministic first-order methods. Here an additional reduction is required because a multiobjective first-order oracle returns strictly richer information than a scalar oracle: at each query $x$ it provides the full collection $\{(f_{i}(x),\nabla f_{i}(x))\}_{i=1}^{m}$, whereas in the scalar setting one observes only $(g(x_{V}),\nabla g(x_{V}))$. To legitimately invoke such scalar minimax lower bounds, we therefore use a simulation (transcript) argument: for the lifted family, any deterministic multiobjective algorithm can be simulated by a deterministic scalar first-order method on $g$ with the same number of oracle calls (since the multiobjective replies can be reconstructed from $(g(x_{V}),\nabla g(x_{V}))$ and the known quadratic $W$-terms). Consequently, if a multiobjective method were able to beat the scalar value-gap lower bound on the lifted instances, the simulated scalar method would also beat it, yielding a contradiction.
Theorem 4.10 (Geometric convex lower bound for adaptive methods under a bounded scalarization gap).
Fix $L>0$, $\Delta>0$, $T\in\mathbb{N}$, and $m\geq 2$. Consider the standard vector first-order oracle for MOO that, at any query $x\in\mathbb{R}^{d}$, returns the transcript item
$\mathcal{O}_{F}(x)\;:=\;\bigl((f_{i}(x),\nabla f_{i}(x))\bigr)_{i=1}^{m}.$
Let $\mathsf{Alg}$ be any deterministic first-order method that makes $T$ oracle calls to $\mathcal{O}_{F}$ and outputs iterates $x^{(0)},x^{(1)},\ldots,x^{(T)}$. Then there exists an $m$-objective MOO instance $F=(f_{1},\ldots,f_{m})$ with the following properties:
  1. (i)
    Each $f_{i}$ is convex and $L$-smooth, the objectives are distinct, and the Pareto set $\mathcal{P}$ is non-singleton (in fact, it contains a non-singleton polytope subset).
  2. (ii)
    (Bounded scalarization gap.) For the scalarization $\lambda=e_{1}$,
    $\Delta_{e_{1}}(x^{(0)})\;:=\;f_{e_{1}}(x^{(0)})-\inf_{x}f_{e_{1}}(x)\;=\;f_{1}(x^{(0)})-\inf_{x}f_{1}(x)\;\leq\;\Delta.$
  3. (iii)
    (Min-iterate lower bound.) The iterates satisfy
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
Equivalently, any deterministic first-order method needs $T=\Omega(\sqrt{L\Delta}/\varepsilon)$ oracle calls to guarantee $\min_{t\leq T}\mathcal{G}(x^{(t)})\leq\varepsilon$ on this class of instances.
Proof.
Oracle model as deterministic transcript maps. A deterministic first-order method $\mathsf{Alg}$ interacting with $\mathcal{O}_{F}$ can be formalized as a sequence of deterministic maps
$\Phi_{t}:\Big(\mathbb{R}^{d}\times(\mathbb{R}\times\mathbb{R}^{d})^{m}\Big)^{t}\to\mathbb{R}^{d},\qquad t\geq 0,$
such that, for $t=0,1,\ldots,T-1$,
$x^{(t+1)}=\Phi_{t}\Big(x^{(0)},\mathcal{O}_{F}(x^{(0)}),\ldots,x^{(t)},\mathcal{O}_{F}(x^{(t)})\Big).$
This is the standard deterministic first-order oracle model; $\mathsf{Alg}$ may use all past queried points and oracle replies.
Step 1: Choose a scalar hard instance from Carmon et al. Apply Lemma 4.3 to the deterministic first-order method that will be constructed below (it will make $T$ scalar first-order oracle calls). Thus there exists a convex $L$-smooth function $g:V\to\mathbb{R}$ (with $\dim(V)\geq 2T$) and an initialization $u^{(0)}\in V$ satisfying
$g(u^{(0)})-\inf_{u\in V}g(u)\leq\Delta$ (12)
and such that the iterates produced by that scalar method obey
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$ (13)
Step 2: Lift $g$ to a non-degenerate MOO instance. Let $W$ be an orthogonal complement of $V$ with $\dim(W)\geq m-1$ and pick affinely independent points $a_{1},\ldots,a_{m}\in W$. Define $F=(f_{1},\ldots,f_{m})$ on $\mathbb{R}^{d}=V\oplus W$ by the lifting construction (Theorem 4.1) with $\gamma=L$:
$f_{i}(u,w)\;:=\;g(u)+\frac{L}{2}\|w-a_{i}\|^{2},\qquad i=1,\ldots,m.$
By Theorem 4.1(i)–(iv), each $f_{i}$ is convex and $L$-smooth, the objectives are distinct, and the Pareto set is
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\ldots,a_{m}\},$
hence non-singleton. In particular, for any $u^{\star}\in\arg\min g$, the slice $\{u^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}\subseteq\mathcal{P}$ is a non-singleton polytope.
We choose the initialization
$x^{(0)}:=(u^{(0)},a_{1}).$
Step 3: Verify the bounded scalarization gap $\Delta_{e_{1}}(x^{(0)})\leq\Delta$. For $\lambda=e_{1}$, $f_{e_{1}}=f_{1}$ and
$f_{1}(u,w)=g(u)+\frac{L}{2}\|w-a_{1}\|^{2}.$
Hence $\inf_{(u,w)}f_{1}(u,w)=\inf_{u}g(u)$ (attained at $(u^{\star},a_{1})$), and therefore
$\Delta_{e_{1}}(x^{(0)})=f_{1}(u^{(0)},a_{1})-\inf_{(u,w)}f_{1}(u,w)=g(u^{(0)})-\inf_{u}g(u)\leq\Delta,$
by (12). This proves item (ii). Importantly, this does not make the MOO instance degenerate: for $i\neq 1$, the objectives differ because $a_{i}\neq a_{1}$. The role of $e_{1}$ is only to express a scalar value-gap condition needed to invoke Lemma 4.3.
Step 4: A transcript-level reduction from MOO to the scalar oracle for $g$. Let $\mathcal{O}_{g}$ be the scalar first-order oracle for $g$, returning $\mathcal{O}_{g}(u):=(g(u),\nabla g(u))$. We now define a deterministic scalar first-order method $\widetilde{\mathsf{Alg}}$ interacting with $\mathcal{O}_{g}$ that simulates the transcript of $\mathsf{Alg}$ interacting with $\mathcal{O}_{F}$ on the lifted instance.
Inductively suppose $\widetilde{\mathsf{Alg}}$ has constructed points $(u^{(0)},w^{(0)}),\ldots,(u^{(t)},w^{(t)})$ equal to the iterates of $\mathsf{Alg}$ on $F$. It queries $\mathcal{O}_{g}$ at $u^{(t)}$ and receives $(g(u^{(t)}),\nabla g(u^{(t)}))$. Using the known lift parameters $L$ and $(a_{i})_{i=1}^{m}$, it reconstructs the full vector-oracle reply at $x^{(t)}=(u^{(t)},w^{(t)})$ as
$\mathcal{O}_{F}(x^{(t)})=\Bigl(g(u^{(t)})+\tfrac{L}{2}\|w^{(t)}-a_{i}\|^{2},\ (\nabla g(u^{(t)}),\,L(w^{(t)}-a_{i}))\Bigr)_{i=1}^{m}.$
Since $\mathsf{Alg}$ is deterministic and its next iterate is given by the deterministic map $\Phi_{t}$ applied to the transcript, $\widetilde{\mathsf{Alg}}$ can compute
$x^{(t+1)}=\Phi_{t}\Big(x^{(0)},\mathcal{O}_{F}(x^{(0)}),\ldots,x^{(t)},\mathcal{O}_{F}(x^{(t)})\Big)$
and set $(u^{(t+1)},w^{(t+1)})=x^{(t+1)}$. Thus the sequence $(u^{(t)})_{t=0}^{T}$ generated by $\widetilde{\mathsf{Alg}}$ is exactly the $V$-projection of the iterates generated by $\mathsf{Alg}$ on the lifted instance. Moreover, $\widetilde{\mathsf{Alg}}$ makes exactly one scalar oracle call per iteration, hence $T$ calls total.
Therefore, the Carmon lower bound (13) applies to the simulated sequence $(u^{(t)})$.
Step 5: Transfer to the Pareto stationarity gap. By Theorem 4.1(iii), for all $t$,
$\mathcal{G}(x^{(t)})\;\geq\;\|\nabla g(u^{(t)})\|.$
Taking the minimum over $0\leq t\leq T$ and using (13) gives
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
This proves item (iii) and completes the proof. ∎
Theorem 4.10 is naturally stated in terms of a scalarization value gap $\Delta$. On our lifted family one can calibrate $\Delta$ and the Pareto distance scale $R$ simultaneously, yielding the $LR/T$ form.
Corollary 4.11 (From $\Delta$ to $R$ on the lifted family).
Fix $L>0$, $R>0$, $T\in\mathbb{N}$, and $m\geq 2$. There exists an MOO instance with $m$ distinct convex $L$-smooth objectives and an initialization $x^{(0)}$ such that:
  1. (i)
    $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ and $\mathcal{P}$ is non-singleton (in particular, it contains a non-singleton polytope subset);
  2. (ii)
    for $\lambda=e_{1}$ the scalarization gap satisfies $\Delta_{e_{1}}(x^{(0)})\asymp LR^{2}$ (up to universal constants);
  3. (iii)
    for any deterministic first-order method run for $T$ steps,
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;c_{0}\,\frac{LR}{T+1}$
    for a universal constant $c_{0}>0$.
Proof.
Let $(g_{0},u_{0}^{(0)})$ be the base hard instance from Theorem 4.10 for parameters $(L_{0},\Delta_{0})$, and denote $u_{0}^{\star}\in\arg\min g_{0}$ and $R_{0}:=\|u_{0}^{(0)}-u_{0}^{\star}\|>0$. For $a,s>0$, define the scaled function $g(u):=a\,g_{0}(su)$ and initialization $u^{(0)}:=u_{0}^{(0)}/s$. Then $\nabla g$ is $L=as^{2}L_{0}$-Lipschitz and
$g(u^{(0)})-\inf g=a\bigl(g_{0}(u_{0}^{(0)})-\inf g_{0}\bigr)=a\Delta_{0}=:\Delta.$
Moreover, since $\arg\min g=\{u_{0}^{\star}/s:u_{0}^{\star}\in\arg\min g_{0}\}$, we have
$\mathrm{dist}(u^{(0)},\arg\min g)=\frac{1}{s}\,\mathrm{dist}(u_{0}^{(0)},\arg\min g_{0})=\frac{R_{0}}{s}.$
Choose $s:=R_{0}/R$ and $a:=L/(s^{2}L_{0})$. Then $\nabla g$ is $L$-Lipschitz and
$\Delta=a\Delta_{0}=\frac{L\Delta_{0}}{s^{2}L_{0}}=\frac{\Delta_{0}}{L_{0}R_{0}^{2}}\,LR^{2}\asymp LR^{2},$
where the implied constants depend only on the fixed base instance.
Now lift $g$ using Theorem 4.1 with $m\geq 2$ distinct points $a_{1},\dots,a_{m}$ in $W$ and parameter $\gamma=L$, and set the initialization $x^{(0)}:=(u^{(0)},a_{1})$. By Theorem 4.1, the Pareto set is $\mathcal{P}=\arg\min g\times\mathrm{conv}\{a_{i}\}$, hence it is non-singleton, and since $(u^{\star},a_{1})\in\mathcal{P}$ for any $u^{\star}\in\arg\min g$ we have
$\mathrm{dist}(x^{(0)},\mathcal{P})=\mathrm{dist}(u^{(0)},\arg\min g)=R.$
Also, for $\lambda=e_{1}$ we have $f_{\lambda}=f_{1}(u,w)=g(u)+\frac{L}{2}\|w-a_{1}\|^{2}$, so $\Delta_{e_{1}}(x^{(0)})=g(u^{(0)})-\inf g=\Delta\asymp LR^{2}$.
Finally, applying Theorem 4.10 yields
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\geq\frac{\sqrt{L\Delta}}{4(T+1)}\asymp\frac{LR}{T+1},$
which proves the claim for a universal constant $c_{0}>0$. ∎
4.5 Nonconvex MOO with Lipschitz gradients
We now use the convexity-free nature of Theorem 4.1 to transfer classical nonconvex stationary-point lower bounds to Pareto stationarity. We emphasize that in nonconvex MOO, $\mathcal{G}(x)=0$ is a necessary stationarity condition (Lemma 3.6) but generally not sufficient for Pareto optimality.
Theorem 4.12 (Nonconvex MOO lower bound under $L$-Lipschitz gradients).
There exists a universal constant $c>0$ such that for any $L>0$, $\Delta>0$, $T\in\mathbb{N}$, and $m\geq 2$, the following holds. For any deterministic first-order method making $T$ vector-oracle calls, there exists an MOO instance with $m$ distinct objectives, each $L$-smooth (not necessarily convex), and an initialization $x^{(0)}$ such that:
  1. (i)
    For the scalarization $\lambda=e_{1}$, $\Delta_{e_{1}}(x^{(0)})\leq\Delta$.
  2. (ii)
    The iterates satisfy
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;c\,\sqrt{\frac{L\Delta}{T+1}}.$
Equivalently, any deterministic first-order method requires $T=\Omega(L\Delta/\varepsilon^{2})$ oracle calls to guarantee $\min_{t\leq T}\mathcal{G}(x^{(t)})\leq\varepsilon$.
Proof.
Apply Lemma 4.4 to obtain an $L$-smooth (possibly nonconvex) scalar function $g:V\to\mathbb{R}$ with $g(u^{(0)})-\inf g\leq\Delta$ but
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\geq c\sqrt{\frac{L\Delta}{T+1}}$
for any deterministic first-order method run for $T$ steps.
Lift $g$ to an MOO instance via Theorem 4.1 with $\gamma=L$ and initialization $x^{(0)}=(u^{(0)},w^{(0)})$ where $w^{(0)}\in\mathrm{conv}\{a_{i}\}$. As in the convex proof, $\Delta_{e_{1}}(x^{(0)})=g(u^{(0)})-\inf g\leq\Delta$. Moreover, by (7),
$\mathcal{G}(x^{(t)})\geq\|\nabla g(u^{(t)})\|\quad\forall t.$
Finally, the same simulation argument as in Theorem 4.10 shows that the $V$-coordinates of any MOO first-order method on the lifted instance correspond to a valid scalar first-order method on $g$, so the scalar lower bound transfers:
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\geq\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\geq c\sqrt{\frac{L\Delta}{T+1}}.$
5 Upper Bounds via Scalarization and Rate Comparison
The lower bounds can be matched by applying an optimal single-objective first-order method to a fixed scalarization of the multiobjective problem. Throughout this section, for any fixed $\lambda\in\Delta^{m}$ we consider the scalarized objective $f_{\lambda}(x):=\sum_{i=1}^{m}\lambda_{i}f_{i}(x)$, which is $L$-smooth (and $\mu$-strongly convex when stated). We write $x_{\lambda}^{\star}\in\arg\min f_{\lambda}$ and $R_{\lambda}:=\|x^{(0)}-x_{\lambda}^{\star}\|$.
To establish upper bounds that match the orders of our lower bounds, we now analyze the performance of an optimal first-order method applied to a fixed scalarization $f_{\lambda}$. We recall the celebrated accelerated gradient method (AGD) of Nesterov. The convergence guarantees for AGD are typically stated in terms of the objective function suboptimality, $f(x)-f^{*}$. Lemma 5.1 provides the crucial link needed to translate these guarantees into bounds on the gradient norm.
Lemma 5.1 (Descent lemma for the gradient norm).
If $f$ is an $L$-smooth convex function with minimum value $f^{\star}$, then for all $x$,
$\|\nabla f(x)\|^{2}\;\leq\;2L\big(f(x)-f^{\star}\big).$
Definition 5.2 (AGD on a fixed scalarization).
Let $f_{\lambda}$ be $L$-smooth (and $\mu$-strongly convex when stated). We use the standard Nesterov accelerated gradient method:
  • Convex case ($\mu=0$). Initialize $x^{0}=y^{0}\in\mathbb{R}^{d}$ and $t_{0}=1$. For $k=0,1,\dots,T-1$,
    $x^{k+1}$ $=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),$
    $t_{k+1}$ $=\tfrac{1+\sqrt{1+4t_{k}^{\,2}}}{2},$
    $y^{k+1}$ $=x^{k+1}+\tfrac{t_{k}-1}{t_{k+1}}\bigl(x^{k+1}-x^{k}\bigr).$
  • Strongly convex case ($\mu>0$). Let $q:=\sqrt{\mu/L}$ and $\beta:=\tfrac{1-q}{1+q}$. Initialize $x^{0}=y^{0}\in\mathbb{R}^{d}$. For $k=0,1,\dots,T-1$,
    $x^{k+1}$ $=y^{k}-\tfrac{1}{L}\,\nabla f_{\lambda}(y^{k}),$
    $y^{k+1}$ $=x^{k+1}+\beta\bigl(x^{k+1}-x^{k}\bigr).$
The convergence properties of this algorithm are well-established and provide optimal rates for the class of first-order methods on smooth convex and smooth, strongly convex functions. We summarize these classical results in the following lemma for completeness.
Lemma 5.3 (AGD convergence (see [28], §2.2; [5], §3.7.2; [4]).
Let $f$ be $L$-smooth and convex with minimizer $x^{\star}$ and $R:=\|x^{(0)}-x^{\star}\|$. The AGD iterates from Definition 5.2 satisfy
$f\big(x^{(T)}\big)-f(x^{\star})\;\leq\;\frac{2L\,R^{2}}{(T+1)^{2}}.$
If, in addition, $f$ is $\mu$-strongly convex with condition number $\kappa:=L/\mu$, then
$f\big(x^{(T)}\big)-f(x^{\star})\;\leq\;\frac{L+\mu}{2}\,R^{2}\,\Bigg(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg)^{\!2T}.$
[…]
Theorem 5.4 (Upper bounds via AGD on a fixed scalarization).
Assume each $f_{i}$ is convex and $L$-smooth. Fix any $\lambda\in\Delta^{m}$ and run AGD (Definition 5.2) on $f_{\lambda}$ for $T$ steps from $x^{(0)}$. Then
$\mathcal{G}\big(x^{(T)}\big)\;\leq\;\|\nabla f_{\lambda}(x^{(T)})\|\;\leq\;\frac{2LR_{\lambda}}{T+1}.$
If, moreover, each $f_{i}$ is $\mu$-strongly convex (so $f_{\lambda}$ is $\mu$-strongly convex with $\kappa=L/\mu$), then
$\mathcal{G}\big(x^{(T)}\big)\;\leq\;\|\nabla f_{\lambda}(x^{(T)})\|\;\leq\;\sqrt{\,L(L+\mu)\,}\;R_{\lambda}\;\Bigg(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg)^{\!T}.$
Proof.
For any $x$, $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$ by definition of the Pareto stationarity gap. In the convex case, combine Lemma 5.1 with Lemma 5.3 applied to $f_{\lambda}$:
$\|\nabla f_{\lambda}(x^{(T)})\|^{2}\;\leq\;2L\big(f_{\lambda}(x^{(T)})-f_{\lambda}(x_{\lambda}^{\star})\big)\;\leq\;2L\cdot\frac{2LR_{\lambda}^{2}}{(T+1)^{2}}\;=\;\frac{4L^{2}R_{\lambda}^{2}}{(T+1)^{2}},$
and take square roots. In the strongly convex case, apply the same inequality $\|\nabla f_{\lambda}(x)\|^{2}\leq 2L(f_{\lambda}(x)-f_{\lambda}^{\star})$ together with the strongly convex part of Lemma 5.3:
$\|\nabla f_{\lambda}(x^{(T)})\|^{2}\;\leq\;2L\cdot\frac{L+\mu}{2}\,R_{\lambda}^{2}\,\Bigg(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg)^{\!2T}\;=\;L(L+\mu)\,R_{\lambda}^{2}\,\Bigg(\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1}\Bigg)^{\!2T},$
and take square roots to obtain the stated bound. Note that AGD is an oblivious span method (but not an oblivious one-step method). Hence an $\mathcal{O}(1/T)$ upper bound in the convex case is achievable by an oblivious span method on a fixed scalarization; the linear bound in the strongly convex case follows analogously with the classical accelerated factor. ∎
Corollary 5.5 (AGD $\varepsilon$-complexity for the Pareto gap).
Under the assumptions of Theorem 5.4 (convex case), to guarantee $\mathcal{G}(x^{(T)})\leq\varepsilon$ it suffices to take
[…]
In the $\mu$-strongly convex case with $\kappa=L/\mu$, it suffices to take
[…]
that is, $T=\Theta\!\big(\sqrt{\kappa}\,\log(\sqrt{L(L+\mu)}\,R_{\lambda}/\varepsilon)\big)$.
Remark 5.6 (Tighter scalar upper bounds via gradient-minimization methods).
The bounds in Theorem 5.4–Corollary 5.5 are obtained by applying AGD to a fixed scalarization $f_{\lambda}$ and then translating function-value convergence into a gradient-norm bound via Lemma 5.1 (using that $G(x)\leq\|\nabla f_{\lambda}(x)\|$). If one is specifically interested in gradient-norm complexity for the scalar problem $\min_{x}f_{\lambda}(x)$, then tighter results are available in the single-objective literature.
In particular, Lan, Ouyang and Zhang [22] develop optimal (and parameter-free) first-order methods for computing $\hat{x}$ with small (projected) gradient norm. Specializing to our unconstrained setting (where projected gradient reduces to the gradient), their results imply that for convex $L$-smooth $f_{\lambda}$ one can obtain $\|\nabla f_{\lambda}(\hat{x})\|\leq\varepsilon$ within $\mathcal{O}\!\left(\sqrt{L}\,R_{\lambda}/\varepsilon\right)$ gradient evaluations, where $R_{\lambda}=\|x^{(0)}-x^{\star}_{\lambda}\|$. In the $\mu$-strongly convex case, they obtain $\mathcal{O}\!\left(\sqrt{L/\mu}\,\log(\|\nabla f_{\lambda}(x^{(0)})\|/\varepsilon)\right)$ gradient evaluations. Combining these guarantees with $G(x)\leq\|\nabla f_{\lambda}(x)\|$ yields the corresponding Pareto-gap upper bounds for the multiobjective problem when optimizing a fixed scalarization.
Remark 5.7 (Best known one-step upper bound).
For $L$-smooth convex $f_{\lambda}$ and stepsizes $\alpha_{t}\leq 1/L$, gradient descent satisfies
$\sum_{t=0}^{T-1}\|\nabla f_{\lambda}(x^{(t)})\|^{2}\;\leq\;2L\big(f_{\lambda}(x^{(0)})-f_{\lambda}^{\star}\big)\;\leq\;L^{2}R_{\lambda}^{2},$
whence $\min_{0\leq t<T}\|\nabla f_{\lambda}(x^{(t)})\|\leq LR_{\lambda}/\sqrt{T}$. Since $\mathcal{G}(x)\leq\|\nabla f_{\lambda}(x)\|$, we obtain $\min_{0\leq t<T}\mathcal{G}(x^{(t)})\leq LR_{\lambda}/\sqrt{T}$ for the oblivious one-step class.
Remark 5.8 (Instance calibration and rate comparison).
The upper bounds above depend on $R_{\lambda}=\|x^{(0)}-x_{\lambda}^{\star}\|$, whereas our lower bounds depend on $R=\mathrm{dist}(x^{(0)},\mathcal{P})$. Always $R\leq R_{\lambda}$. For the hard instances in Theorem 4.1, choosing $\lambda=e_{1}$ yields $R_{e_{1}}=R$, so the orders match. Importantly, the $\Omega(1/T)$ lower bound in Theorem 4.6 holds for the oblivious one-step class, whereas the $\mathcal{O}(1/T)$ upper bound here is achieved by an oblivious span method (AGD). Whether $\mathcal{O}(1/T)$ is achievable within the oblivious one-step class remains open; the best known upper bound there is $\mathcal{O}(LR/\sqrt{T})$.
6 Conclusion
We developed an oracle-complexity theory for finding $\varepsilon$-Pareto stationary points in smooth multiobjective optimization, measured by the Pareto stationarity gap $\mathcal{G}(x)$. A central tool is a robust non-degenerate lifting construction that embeds scalar hard instances into MOO instances with distinct objectives and non-singleton Pareto fronts while preserving lower bounds on $\mathcal{G}$.
In the $\mu$-strongly convex regime we obtained a tight linear-rate lower bound for span methods matching accelerated upper bounds. In the convex regime we separated algorithmic power: oblivious one-step methods admit an $\Omega(1/T)$ min-iterate lower bound, while oblivious span methods satisfy a universal polynomial-degree $\Omega(1/T^{2})$ last-iterate bound. We then showed that the polynomial-degree limitation is loose: by lifting the geometric stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford, we proved an $\Omega(\sqrt{L\Delta}/T)$ min-iterate lower bound for adaptive first-order methods under bounded scalarization gap, recovering $\Omega(LR/T)$ on our lifted family. Finally, leveraging that the lifting does not require convexity of the embedded function, we extended the framework to nonconvex MOO under Lipschitz gradients, obtaining the tight-in-order $\Omega(\sqrt{L\Delta/T})$ complexity.
Several directions remain open. Most prominently, the last-iterate complexity of adaptive first-order methods in convex MOO under bounded scalarization gap (or under geometric distance models) is not yet fully characterized. Extending the framework to randomized methods, stochastic oracles, constraints/composite objectives, and refined dimension-dependent regimes are further natural directions for a comprehensive complexity theory of Pareto stationarity.
Appendix A Derivations of Polynomial Extremal Results
A.1 A tight $\Omega(1/T)$ lower bound for oblivious one-step methods (with $L$-capped steps)
In this appendix we establish a tight $\Omega(1/T)$ lower bound for the stationarity-gap proxy
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr),$
which arises in the analysis of oblivious one-step methods on $L$-smooth convex quadratics. The proof is self-contained and uses only elementary inequalities.
Model and class
We consider oblivious one-step gradient methods applied to a fixed scalarization $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ of a multiobjective instance, where $H\succeq 0$ and $\sigma(H)\subset[0,L]$. The method generates
$x^{(t+1)}\;=\;x^{(t)}-\alpha_{t}\,\nabla g\bigl(x^{(t)}\bigr),\qquad t=0,1,2,\dots,$
with a pre-scheduled stepsize sequence $\{\alpha_{t}\}_{t\geq 0}$ which may depend on known problem data (here, $L$) but not on oracle feedback. We adopt the natural $L$-cap:
$0\;\leq\;\alpha_{t}\;\leq\;\frac{1}{L}\qquad\text{for all }t\geq 0.$ (14)
This is the standard safe region for gradient descent on $L$-smooth convex quadratics with known $L$ and is the relevant oblivious regime for worst-case analysis.
Polynomial representation
Let $e^{(t)}=x^{(t)}-x^{\star}$, where $x^{\star}\in\arg\min g$; then for quadratics $e^{(t)}=p_{t}(H)\,e^{(0)}$ with
$p_{t}(\zeta)\;=\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr),\qquad p_{0}\equiv 1,$
and hence $\nabla g\bigl(x^{(t)}\bigr)=H\,p_{t}(H)\,e^{(0)}$ (Lemma 3.15).
We prove the following extremal result.
Theorem A.1.1 (Tight $\Omega(1/T)$ product-form extremal under $L$-capped steps).
Fix $L>0$ and $t\in\mathbb{N}$. For any nonnegative stepsizes $\alpha_{0},\ldots,\alpha_{t-1}$ satisfying (14), we have
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr)\;\;\geq\;\;\frac{L}{4\,(t+1)}.$ (15)
Moreover, the bound is tight up to a universal constant: the constant stepsizes $\alpha_{k}\equiv 1/L$ yield
$\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr)\;=\;\max_{\zeta\in[0,L]}\;\zeta\Bigl(1-\frac{\zeta}{L}\Bigr)^{t}\;\;\leq\;\;\frac{L}{e\,(t+1)}.$ (16)
Remark A.1.2 (Scope of (14)).
The cap (14) is standard in first-order analysis with known $L$. Allowing stepsizes larger than $1/L$ may cause ascent on $L$-smooth convex quadratics and does not help the algorithm in the worst case. Our lower bound therefore covers the canonical oblivious regime.
The proof rests on a simple but sharp product inequality.
Lemma A.1.3 (A product lower bound).
For any $r\in\mathbb{N}$ and any numbers $x_{1},\ldots,x_{r}\in[0,1]$,
$\prod_{i=1}^{r}\bigl(1-x_{i}\bigr)\;\;\geq\;\;1-\sum_{i=1}^{r}x_{i}.$ (17)
Proof.
By induction on $r$. The case $r=1$ is trivial. Suppose (17) holds for r1r-1. Then for $r$ we have
$\prod_{i=1}^{r}(1-x_{i})$ $=\Bigl(\prod_{i=1}^{r-1}(1-x_{i})\Bigr)(1-x_{r})$
$\geq\Bigl(1-\sum_{i=1}^{r-1}x_{i}\Bigr)(1-x_{r})$
$=1-\sum_{i=1}^{r}x_{i}\;+\;x_{r}\sum_{i=1}^{r-1}x_{i}$
$\geq 1-\sum_{i=1}^{r}x_{i}.$
since $x_{r}\sum_{i=1}^{r-1}x_{i}\geq 0$. ∎
Proof of Theorem A.1.1.
Fix $t\geq 1$ and a stepsize sequence $(\alpha_{k})_{k=0}^{t-1}$ with $0\leq\alpha_{k}\leq\tfrac{1}{L}$ for all $k$. For any $\zeta\in[0,L]$, set $x_{k}:=\alpha_{k}\,\zeta\in[0,1]$. Applying Lemma A.1.3 gives
$\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr)\;\;\geq\;\;1-\zeta\sum_{k=0}^{t-1}\alpha_{k}.$ (18)
Define the partial sums $S_{t}:=\sum_{k=0}^{t-1}\alpha_{k}$. Note that $S_{t}\leq t\cdot(1/L)=t/L$ by (14).
Consider the function $\Phi_{t}(\zeta):=\zeta\prod_{k=0}^{t-1}(1-\alpha_{k}\,\zeta)$ on $[0,L]$. By (18),
$\Phi_{t}(\zeta)\;\geq\;\zeta\bigl(1-\zeta S_{t}\bigr)\qquad\text{for all }\zeta\in[0,L].$ (19)
Let $\zeta^{\star}:=\min\{L,\;1/(2S_{t})\}$ (with the convention $1/0=+\infty$). We distinguish two exhaustive cases.
Case 1: $S_{t}\geq 1/(2L)$. Then $\zeta^{\star}=1/(2S_{t})\leq L$, and hence, by (19),
$\Phi_{t}(\zeta^{\star})\;\geq\;\zeta^{\star}\Bigl(1-\zeta^{\star}S_{t}\Bigr)\;=\;\frac{1}{2S_{t}}\cdot\Bigl(1-\tfrac{1}{2}\Bigr)\;=\;\frac{1}{4S_{t}}\;\geq\;\frac{1}{4\,(t/L)}\;=\;\frac{L}{4t}.$
Case 2: $S_{t}<1/(2L)$. Then $\zeta^{\star}=L$ and, by (19),
$\Phi_{t}(\zeta^{\star})\;\geq\;L\bigl(1-LS_{t}\bigr)\;\geq\;L\Bigl(1-\tfrac{1}{2}\Bigr)\;=\;\frac{L}{2}\;\geq\;\frac{L}{4t}\qquad(\text{since }t\geq 1).$
In both cases, we obtain $\max_{\zeta\in[0,L]}\Phi_{t}(\zeta)\geq L/(4t)$. Since $t\geq 1$ implies $1/t\geq 1/(t+1)$, this establishes the slightly looser bound in (15).
For the near-matching upper bound (16), take the constant stepsizes $\alpha_{k}\equiv 1/L$. Then
$\Psi_{t}(\zeta):=\zeta\Bigl(1-\frac{\zeta}{L}\Bigr)^{t},\qquad\zeta\in[0,L].$
Elementary calculus shows that $\Psi_{t}$ attains its maximum at $\zeta^{\diamond}=L/(t+1)$, with
$\max_{\zeta\in[0,L]}\Psi_{t}(\zeta)$ $=\Psi_{t}(\zeta^{\diamond})$
$=\frac{L}{t+1}\left(1-\frac{1}{t+1}\right)^{t}$
$\leq\frac{L}{t+1}\,\mathrm{e}^{-\,t/(t+1)}$
$\leq\frac{L}{\mathrm{e}\,(t+1)}.$
This establishes (16). ∎
We now translate Theorem A.1.1 into an oracle lower bound for the gradient norm.
Corollary A.1.4 (Oblivious one-step lower bound for the gradient norm).
Let $t\in\mathbb{N}$ and let an oblivious one-step method with stepsizes $\alpha_{0},\ldots,\alpha_{t-1}$ obeying (14) be given. Then there exists an $L$-smooth convex quadratic $g(x)=\tfrac{1}{2}x^{\top}Hx-b^{\top}x$ with $\sigma(H)\subset[0,L]$ and an initialization $x^{(0)}$ satisfying $\|x^{(0)}-x^{\star}\|=R$ such that
$\min_{0\leq s\leq t}\;\bigl\|\nabla g\bigl(x^{(s)}\bigr)\bigr\|\;\;\geq\;\;\frac{L\,R}{4\,(t+1)}.$
Proof.
For quadratics, $\|\nabla g(x^{(s)})\|=\|H\,p_{s}(H)\,(x^{(0)}-x^{\star})\|$. Fix any unit vector $u$ and take $x^{(0)}-x^{\star}=R\,u$. For every fixed $\zeta\in[0,L]$, the monotonicity $(1-\alpha_{s}\zeta)\in[0,1]$ implies
$\zeta\,\Bigl|\prod_{k=0}^{s-1}\bigl(1-\alpha_{k}\zeta\bigr)\Bigr|\;\;\text{is nonincreasing in $s$.}$
Hence, for any fixed $\zeta$,
$\min_{0\leq s\leq t}\ \zeta\,\Bigl|\prod_{k=0}^{s-1}\bigl(1-\alpha_{k}\zeta\bigr)\Bigr|\;=\;\zeta\,\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\zeta\bigr).$
Choosing $H$ to be diagonal with an eigenvalue at the maximizer $\zeta^{\star}\in[0,L]$ of the right-hand side and aligning $u$ with the corresponding eigenvector yields
$\min_{0\leq s\leq t}\;\|\nabla g(x^{(s)})\|\;=\;R\cdot\max_{\zeta\in[0,L]}\;\zeta\;\prod_{k=0}^{t-1}\bigl(1-\alpha_{k}\,\zeta\bigr)\;\geq\;\frac{L\,R}{4\,(t+1)},$
by Theorem A.1.1. ∎
This result is lifted to the multiobjective setting using Theorem 4.1: for the instance $f_{i}(x)=g(x_{V})+\frac{L}{2}\|x_{W}-a_{i}\|^{2}$, we have $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$. Corollary A.1.4 thus directly implies a lower bound on $\min_{s\leq t}\mathcal{G}(x^{(s)})$. This establishes the bound for a non-degenerate MOO instance with distinct objectives and a non-singleton Pareto set.
A.2 Universal Markov-Based Bound for Oblivious Span Methods
The lower bound for oblivious methods relies on $\inf_{p_{t}}\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|$ where $p_{t}$ is any polynomial with $\deg p_{t}\leq t$ and $p_{t}(0)=1$.
Lemma A.2.1 (Markov Brothers’ Inequality (see [32], Chapter 2, Section 2.7, Page 123)).
For any polynomial $q(x)$ of degree $n$,
$\max_{x\in[-1,1]}|q^{\prime}(x)|\leq n^{2}\max_{x\in[-1,1]}|q(x)|.$
Let $S(\zeta)=\zeta p_{t}(\zeta)$, which is a polynomial of degree $t+1$. We know $S(0)=0$ and from $p_{t}(0)=1$, we have $S^{\prime}(0)=p_{t}(0)+0\cdot p_{t}^{\prime}(0)=1$. Let’s map the interval $[0,L]$ to [1,1][-1,1] via the transformation $x=\frac{2\zeta}{L}-1$, which means $\zeta=\frac{L(x+1)}{2}$. Define a new polynomial $q(x)=S\left(\frac{L(x+1)}{2}\right)$. The degree of $q$ is $n=t+1$. At x=1x=-1, q(1)=S(0)=0q(-1)=S(0)=0. The derivative of $q$ with respect to $x$ is $q^{\prime}(x)=S^{\prime}\left(\frac{L(x+1)}{2}\right)\cdot\frac{L}{2}$. At x=1x=-1, this gives $q^{\prime}(-1)=S^{\prime}(0)\cdot\frac{L}{2}=1\cdot\frac{L}{2}=\frac{L}{2}$. By Markov’s inequality:
$\frac{L}{2}=|q^{\prime}(-1)|\leq\max_{x\in[-1,1]}|q^{\prime}(x)|\leq(t+1)^{2}\max_{x\in[-1,1]}|q(x)|.$
Rearranging gives the bound on the maximum of $q$, which is the same as the maximum of $S$:
$\max_{\zeta\in[0,L]}|\zeta p_{t}(\zeta)|=\max_{x\in[-1,1]}|q(x)|\geq\frac{L}{2(t+1)^{2}}.$
A.3 Chebyshev Extremal Problem on $\mu,L]$
The lower bound for strongly convex methods requires solving
$\min_{p_{t}}\;\max_{\zeta\in[\mu,L]}\lvert p_{t}(\zeta)\rvert\quad\text{s.t.}\quad\deg p_{t}\leq t,\;p_{t}(0)=1.$
The affine transformation $\xi=\frac{2\zeta-(L+\mu)}{L-\mu}$ maps the interval $\zeta\in[\mu,L]$ to $\xi\in[-1,1]$. The constraint $p_{t}(0)=1$ is evaluated at $\zeta=0$, which corresponds to $\xi_{0}=-\frac{L+\mu}{L-\mu}$. Let $\tilde{p}_{t}(\xi)=p_{t}(\zeta(\xi))$. The problem becomes:
$\min_{\begin{subarray}{c}\deg\tilde{p}_{t}\leq t\\ \tilde{p}_{t}(\xi_{0})=1\end{subarray}}\max_{\xi\in[-1,1]}|\tilde{p}_{t}(\xi)|.$
The solution to this classic problem is the scaled Chebyshev polynomial $\tilde{p}_{t}^{*}(\xi)=\frac{T_{t}(\xi)}{T_{t}(\xi_{0})}$, where $T_{t}$ is the Chebyshev polynomial of the first kind of degree $t$ (see [32], Chapter 2, Section 2.7, Page 108). The extremal value is $\frac{1}{|T_{t}(\xi_{0})|}$ because $\max_{\xi\in[-1,1]}|T_{t}(\xi)|=1$. With $\kappa=L/\mu$, we have $|\xi_{0}|=\frac{L+\mu}{L-\mu}=\frac{\kappa+1}{\kappa-1}>1$. For $|x|>1$, the Chebyshev polynomial is given by $T_{t}(x)=\frac{1}{2}\left((x+\sqrt{x^{2}-1})^{t}+(x-\sqrt{x^{2}-1})^{-t}\right)$. Let $\rho=|\xi_{0}|+\sqrt{|\xi_{0}|^{2}-1}$. A direct calculation shows:
$\rho$ $=\frac{\kappa+1}{\kappa-1}+\sqrt{\left(\frac{\kappa+1}{\kappa-1}\right)^{2}-1}$
$=\frac{\kappa+1+2\sqrt{\kappa}}{\kappa-1}$
$=\frac{(\sqrt{\kappa}+1)^{2}}{(\sqrt{\kappa}-1)(\sqrt{\kappa}+1)}$
$=\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}.$
Then $|T_{t}(\xi_{0})|=|T_{t}(|\xi_{0}|)|=\frac{1}{2}(\rho^{t}+\rho^{-t})$. The extremal value is $\frac{1}{|T_{t}(\xi_{0})|}=\frac{2}{\rho^{t}+\rho^{-t}}$.